### v1

# DiffPatch: Generating Customizable Adversarial Patches using Diffusion Model

Abstract
Physical adversarial patches printed on clothing can easily allow individuals to evade person detectors. However, most existing adversarial patch generation methods prioritize attack effectiveness over stealthiness, resulting in patches that are aesthetically unpleasing. Although existing methods using generative adversarial networks or diffusion models can produce more natural-looking patches, they often struggle to balance stealthiness with attack effectiveness and lack flexibility for user customization. To address these challenges, we propose a novel diffusion-based customizable patch generation framework termed DiffPatch, specifically tailored for creating naturalistic and customizable adversarial patches. Our approach enables users to utilize a reference image as the source, rather than starting from random noise, and incorporates masks to craft naturalistic patches of various shapes, not limited to squares. To prevent the original semantics from being lost during the diffusion process, we employ Null-text inversion to map random noise samples to a single input image and generate patches through Incomplete Diffusion Optimization (IDO). Notably, while maintaining a natural appearance, our method achieves a comparable attack performance to state-of-the-art non-naturalistic patches when using similarly sized attacks. Using DiffPatch, we have created a physical adversarial T-shirt dataset, AdvPatch-1K, specifically targeting YOLOv5s. This dataset includes over a thousand images across diverse scenarios, validating the effectiveness of our attack in real-world environments. Moreover, it provides a valuable resource for future research.
1 Introduction
Figure 1: Example images from our AdvPatch-1K dataset, showcasing the performance of DiffPatch in various scenarios.
With the rapid advancement of computer vision, object detectors are extensively utilized in real-world scenarios, including fields such as autonomous driving [8] and medical diagnostics [33]. Unfortunately, recent works [7, 47, 15, 45] have demonstrated that deep neural networks are vulnerable to maliciously crafted adversarial examples, resulting in decisions that defy common sense. This poses significant threats to safety-critical applications.
In the field of computer vision, adversarial attacks can be categorized into two distinct groups: digital attacks [17, 7] and physical attacks [14, 27, 32]. Different from digital attacks, which involve introducing small perturbations within a digital domain, physical adversarial attacks operate in the tangible world, aiming to mislead visual systems through modifications to the appearance or positioning of physical objects. The execution of physical adversarial attacks can be accomplished through numerous methods, including the application of adversarial patches onto the surface of an object, alterations in environmental lighting conditions, or the introduction of distracting elements.
For physical-world adversarial patches, previous research [43, 48] has primarily focused on enhancing attack performance, which often requires introducing stronger perturbations that significantly deviate patches from the distribution of natural images, making them appear highly unnatural. While most patch generation methods [46, 23] incorporate a total variation loss to smooth the patches, this fundamental issue remains unresolved, rendering the patches easily identifiable by humans. With the rapid development of image generation techniques [16, 20], naturalistic patch generation methods based on deep image generators have gradually emerged [21, 30]. The patches generated by sate-of-the-art generators possess a more natural appearance, significantly enhancing their stealthiness. However, these methods suffer from limited attack effectiveness compared to unnatural patches. Moreover, these methods have high randomness in the generation process, making it difficult to generate customized patches based on a given image.
In this work, we propose a novel Stable Diffusion [39] based adversarial patch generation method called DiffPatch that is capable of generating a naturalistic and customizable adversarial patch based on a reference image specified by the adversary. Our DiffPatch method works as follows. Firstly, it maps the reference image into the latent space and generates adversarial patches by optimizing the noise latent vectors. It adopts Null-text inversion [34] to achieve a near-perfect reconstruction so as to guarantee that the generated images largely preserve the original semantics. Secondly, it introduces an Incomplete Diffusion Optimization (IDO) strategy to ensure the natural appearance and attack performance of the optimized patches and employs an IoU-Detection loss to accelerate convergence. Furthermore, to generate irregular-shaped patches, it utilizes masks to replace the background and covers the gradients in non-essential areas. This can also help reduce information interference from other regions during the optimization process. Based on DiffPatch, we have created an adversarial patch dataset AdvPatch-1K which consists of over a thousand images, featuring attacks in various physical-world scenarios with customizable patches of diverse styles and shapes. Partial results are shown in Figure 1. Our main contributions are as follows:
  • We propose a novel diffusion-based method called DiffPatch for generating naturalistic adversarial patches based on a reference image. DiffPatch employs Null-text inversion, Incomplete Diffusion Optimization (IDO) , masked customization, and an IoU-Detection loss to generate stylized and customizable adversarial patches.
  • Experimental results show that our DiffPatch achieves the highest attack success rate (ASR), significantly reducing the mAP of various object detectors, rivaling or even surpassing the current state-of-the-art methods of unnatural adversarial patches.
  • We create physical-world adversarial patches by printing them on T-shirts and collect a dataset, AdvPatch-1K, which consists of 1,131 images with annotations. This dataset covers 9 unique adversarial T-shirt designs across various contexts, including both indoor and outdoor settings as well as individual and group photos.
2 Related Work
2.1 Physical Adversarial Attack
With the widespread application of deep learning, researchers began to investigate the security of these models in real-world scenarios. This focus led to the development of the Basic Iterative Method (BIM) [27], which generates adversarial samples that retain their adversarial properties even after being printed. To improve the robustness of physical adversarial samples, the Expectation Over Transformation (EOT) approach [2] was introduced. EOT enhances resistance to real-world transformations by simulating various modifications that images may undergo. The Digital-to-Physical (D2P) method [24] further advances real-world attack effectiveness by learning the transformations between digital and physical images. AdvPatch [5] marks a milestone by demonstrating the creation of robust adversarial patches that function reliably across diverse environmental conditions. Following this, AdvYOLO [43] demonstrated an attack on object detectors by applying adversarial patches to clothing surfaces, allowing wearers to evade detection. Several similar patch attack methods have since been proposed [48, 46, 52, 22], posing a significant threat to object detectors. Additionally, other physical adversarial attack methods have employed accessories like glasses [41], lighting [51], and shadows [50].
2.2 Naturalistic Adversarial Patch
Inspired by the imperceptibility constraints in digital attacks, Universal Physical Camouflage Attacks (UPC) [23] employ the $L_{\infty}$ norm to limit perturbations and maintain the natural appearance of patches. Unlike iterative methods that optimize patches on images, the Naturalistic Adversarial Patch (NAP) [21] utilizes a generative approach. NAP leverages GANs to indirectly optimize the patch by modifying initial latent vectors, resulting in patches with a highly natural appearance. Building on these advancements, several GAN-based generation methods have emerged [13, 28]. Additionally, diffusion model-based methods have gained traction in the vision field [9, 30, 49]. The generative power of diffusion models enables the creation of more naturalistic and effective physical adversarial patches.
Figure 2: An overview of the proposed DiffPatch generation framework.
2.3 Diffusion-based Image Editing
Recently, diffusion models have excelled in the field of image generation, with models like DALL-E2 [36], Imagen [40], and Stable Diffusion [39] largely surpassing GANs in performance [12]. Beyond generation tasks, diffusion-based image editing techniques have also become mainstream. Leveraging the cross-attention mechanism, Prompt to Prompt (P2P) [18] introduces an innovative framework that enables image editing solely through text prompts. To address challenges with poor reconstruction during editing, Null-text inversion [34] achieves near-perfect reconstruction by optimizing the unconditional embedding at each denoising step. Text-driven Blended Diffusion [3] employs masks to target specific regions in an image, allowing for background replacement and consistent edits within defined areas. Additionally, some adversarial attack methods [49, 10] leverage image editing techniques to modify images while preserving their natural appearance, creating adversarial examples that are visually coherent yet misleading to detection models.
3 Proposed Approach
Overview
Figure 2 provides an overview of DiffPatch. It optimizes the Null text embeddings along incomplete diffusion trajectories and maps the image to the latent vectors at intermediate timestamps, ensuring that the patch maintains a natural appearance during updates. It then applies the Incomplete Diffusion Optimization (IDO) guided by an IoU-Detection Loss to optimize the latent vectors, resulting in faster convergence and enhanced attack performance. During optimization, we skip the U-Net gradients to avoid memory overflow and utilize target masks to eliminate semantic interference from irrelevant areas, thereby generating more naturalistic adversarial patches. Next, we will introduce the key techniques used in DiffPatch.
3.1 Image-to-Latent Mapping
Our objective is to achieve a perfect mapping between a given image $\mathcal{I}$ and latent vector $z_{t}$, utilizing the image encoding $z_{0}$ to obtain $z_{t}$, which is essentially the reverse process of DDIM sampling [42]. Based on the assumption that the ODE process can be reversed in the limit of small steps, DDIM inversion can be formulated as:
$z_{t+1}=\sqrt{\frac{\alpha_{t+1}}{\alpha_{t}}}z_{t}+\left(\sqrt{\frac{1}{\alpha_{t+1}}-1}-\sqrt{\frac{1}{\alpha_{t}}-1}\right)\cdot\varepsilon_{\theta}(z_{t},t,\mathcal{C}),$
where $\varepsilon_{\theta}$ is the pre-trained U-Net used to predict noise, $\mathcal{C}$ is the condition embedding of $\mathcal{P}$, and $\alpha_{t}$ is calculated based on the schedule $\beta_{0},\ldots,\beta_{T}\in(0,1)$, i.e., $\alpha_{t}=\prod_{i=1}^{t}(1-\beta_{i})$. Through the aforementioned formula, we can obtain the latent vector $z_{t}$ and utilize it for image reconstruction, yet this does not guarantee perfect reconstruction.
To enhance the quality of generated outputs, Stable Diffusion incorporates a classifier-free guidance mechanism [19] that concurrently engages in unconditional generation and guided generation based on condition embedding $\mathcal{C}$:
$\tilde{\varepsilon}_{\theta}(z_{t},t,\mathcal{C},\phi)=w\cdot\varepsilon_{\theta}(z_{t},t,\mathcal{C})+(1-w)\ \cdot\varepsilon_{\theta}(z_{t},t,\phi),$ (2)
where $w$ represents the guidance scale parameter, and $\phi=\mathcal{E}("")$ is the embedding of a null text. The generated content is highly sensitive to textual prompts. During each step of guided generation, even minor variations in the prompt can lead to deviations in the trajectory. This cumulative error can make reconstruction challenging, and such errors tend to amplify when we perform image editing.
To resolve this issue, we separately optimize the null text embedding $\phi_{t}$ for each timestamp $t$ in the order of the diffusion process $t=T\rightarrow t=1$, using the optimized embedding to guide the process back onto the correct trajectory. Similar to Null-text inversion [34], firstly, we set the guidance scale $w$ to 1 and utilize DDIM inversion to obtain intermediate vectors $z^{*}_{T},\ldots,z^{*}_{0}$ where $z^{*}_{0}=z_{0}$. After that, regarding $z^{*}_{T}$ as the initial noise vector $\bar{z}_{T}$, we optimize the null text embedding $\phi_{t}$ through the application of the following objective function, after each iteration of the DDIM sampling step (adopting the default value of 7.5 for the guidance scale $w$):
$\min_{\phi_{t}}\|z^{*}_{t-1}-z_{t-1}(\bar{z_{t}},\phi_{t},\mathcal{C})\|^{2}_{2},$ (3)
where $z_{t-1}$ represents the vector obtained after DDIM sampling, and after each step of optimization is completed, we update $\bar{z}_{t-1}=z_{t-1}$. Finally, we can achieve near-perfect reconstruction by utilizing the initial noise vector $\bar{z}_{T}$ and the optimized unconditional embeddings $\{\phi_{t}\}^{T}_{t=1}$.
3.2 Incomplete Diffusion Optimization
We propose Incomplete Diffusion Optimization (IDO), which adds adversarial perturbations along the trajectory of image reconstruction. IDO consists of two techniques: 1) IoU-Detection Loss and 2) Incomplete Diffusion Process.
IoU-Detection Loss
Loss function is crucial in guiding the optimization process. After an intermediate adversarial image is fed into the detector, we obtain the following outputs: 1) $B_{box}$: coordinates of the predicted bounding box; 2) $P_{obj}$: probability indicating the presence of an object within the predicted box; and 3) $P_{cls}$: classification probabilities for various objects. Previous work used Common Detection Loss [43] to guide adversarial patch generation:
$\mathcal{L}_{det}=\frac{1}{N}\sum^{N}_{i=1}\max_{\begin{subarray}{c}j\end{subarray}}(P^{j}_{obj}(I^{\prime}_{i})\cdot P^{j}_{cls}(I^{\prime}_{i})),$ (4)
where $N$ represents the batch size, and $I^{\prime}_{i}$ refers to the $i$-th image in a batch. For each image $I^{\prime}_{i}$, the detector identifies multiple objects and assigns an index $j$ to each detected object. The images $\{I^{\prime}_{i}\}_{i=1}^{N}$ are obtained by performing operation $\mathcal{T}$ on clean images, which means pasting patches on the surface of the clothes of the person inside.
According to the loss function formulation in Equation (4), the optimization for a given image $I^{\prime}_{i}$ is guided solely by the object with the highest confidence product, aiming to minimize the detection probability for that specific target. However, this strategy introduces several challenges. First, the application of adversarial patches is constrained by the availability and accuracy of dataset labels. Incomplete annotations or misdetections can result in critical surfaces remaining unpatched, undermining the optimization process. Moreover, focusing the loss on a single object per iteration leads to slower convergence and increased instability during training, limiting the overall effectiveness of the approach.
We propose an alternative loss function IoU-Detection Loss to address these issues. Formally, it is defined as:
$\mathcal{L}_{IoU}=\frac{1}{N}\sum_{i=1}^{N}\left(\frac{1}{M}\sum_{k=1}^{K}\left[\mathbbm{1}\left(\max_{j}\mathrm{IoU}(J_{j},J^{\prime}_{k})>t\right)P(J^{\prime}_{k})\right]\right),$
where $M$ is the total number of detected bounding boxes with an Intersection over Union (IoU) greater than a threshold $t$ with any ground truth box, $\mathrm{IoU}(J_{j},J^{\prime}_{k})$ represents the IoU between the predicted bounding box $J^{\prime}_{k}$ and the ground truth box $J_{j}$, $\mathbbm{1}(\cdot)$ is an indicator function that equals 1 if the condition inside the parentheses is met and 0 otherwise, and $P$ denotes the product of the object probability $P_{obj}$ and the classification probability $P_{cls}$. The IoU-Detection Loss accounts for multiple objects with patches applied to a single image and mitigates the impact of non-patched objects on the training process.
Incomplete Diffusion Process
We found that using the default 50-step sampling to generate patches often leads to a loss of semantic meaning and natural appearance during optimization. Excessive sampling steps can amplify subtle changes in $z_{T}$, causing the image reconstruction trajectory to deviate. During the image-to-latent mapping phase, we can match it with the latent vector $z_{\frac{T}{2}}$ at an intermediate timestamp to mitigate semantic loss. The generation of adversarial patch $z_{p}$ can then be represented as:
$z_{p}=G(\bar{z}_{\frac{T}{2}},\frac{T}{2},\mathcal{C},\{\phi_{t}\}^{\frac{T}{2}}_{t=1}),$ (6)
where $G$ represents the diffusion model. Generating patches using $z_{\frac{T}{2}}$ shortens the generation path, thereby enhancing the stability of the patches under adversarial guidance. However, as $z_{\frac{T}{2}}$ continues to be updated, it becomes more chaotic, making complete denoising challenging in the later stages of the process. To prevent adversarial perturbations from being noticeable, perturbation values are constrained by an infinity norm in digital attacks. Similarly, we constrain the perturbations $\delta$ applied to the latent vector $z_{\frac{T}{2}}$:
$\delta_{t}=\text{Proj}_{\infty}(\delta_{t-1}+\Delta\delta,0,\epsilon),$ (7)
where the perturbation $\delta$ is constrained within a sphere centered at 0 with a radius of $\epsilon$ through the projection function. This ensures that the internal semantics of $z_{\frac{T}{2}}$ are not excessively lost, enabling the generation of patches with natural appearances through the denoising process.
Specifically, during denoising, 25 steps of DDIM sampling are required, where U-Net performs 25 noise predictions. The denoising process involves removing noise at timestamp $t$ using a sampling recurrence formula to obtain the vector $z_{t-1}$, then gradually obtaining $z_{0}$, and finally using an image decoder $\Psi$ to restore the image. When optimizing the vector $z_{T/2}$ with the loss function $\mathcal{L}$, it can be expanded using the chain rule as follows:
$\nabla_{z_{T/2}}\mathcal{L}=\frac{\partial\mathcal{L}}{\partial x}\cdot\frac{\partial x}{\partial z_{p}}\cdot\frac{\partial z_{p}}{\partial z_{0}}\cdot\frac{\partial z_{0}}{\partial z_{1}}\cdot\cdots\cdot\frac{\partial z_{T/2-1}}{\partial z_{T/2}},$ (8)
where $z_{p}$ represents the adversarial patch obtained through the image encoder $\Psi$, i.e., $z_{p}=\Psi(z_{0})$ , and $x$ is the training data with $z_{p}$ attached. For each step of DDIM sampling, it can be expressed as:
[…]
Each denoising step requires computing and storing the U-Net gradient, which can lead to memory overflow. However, since the predicted noise $\varepsilon_{\theta}(z_{t},t,\mathcal{C})$ follows a standard normal distribution $\mathcal{N}(0,I)$, we can use the reparameterization trick [10] to bypass the gradient computation of the U-Net and obtain an approximate gradient:
$\nabla_{z_{T/2}}\mathcal{L}\approx\frac{\partial\mathcal{L}}{\partial x}\cdot\frac{\partial x}{\partial z_{p}}\cdot\frac{\partial z_{p}}{\partial z_{0}}\cdot\sqrt{\frac{1}{\alpha_{1}}}\cdot\cdots\cdot\sqrt{\frac{\alpha_{T/2-1}}{\alpha_{T/2}}}.$ (10)
3.3 Target Mask Control
Our method can generate adversarial patches of various shapes, with the process being controlled by the target mask $m$. Using target masks serves not only to extract patches of different shapes but also plays a crucial role in the natural appearance of the patch. Since Stable Diffusion [39] generates square images, there is environmental information surrounding our target. After selecting the initial image $\mathcal{I}$, to prevent the background from interfering with the main subject during the optimization process, we need to replace the background with a solid-colored image $s$ before the Image-to-Latent Mapping stage:
$\mathcal{I}^{\prime}=\mathcal{I}\odot m+s\odot(1-m),$ (11)
Additionally, to ensure that the background remains unchanged throughout the update process, we downsample the mask $m$ by a factor of 8 to match the size of latent vectors and overlay gradients outside the target area.
Overall, the DiffPatch generation process starts with Image-to-Latent Mapping. Incomplete Diffusion Optimization is then applied to reduce memory consumption and enhance performance. Finally, Target Mask Control is used to maintain a natural appearance. The generation process is illustrated in Algorithm 1. In addition, the patch can be iteratively optimized using DiffPatch to achieve an accumulative effect, as shown in Figure 3. Specifically, the generated patch can be used to generate a new patch, enabling stronger attack performance while maintaining natural appearance.
Figure 3: The effect of iterative generation (accumulated style variation from left to right).
Algorithm 1 DiffPatch Generation Process
Input: A patch image $\mathcal{I}^{\prime}$ with target mask $m$, prompt embedding $\mathcal{C}=\mathcal{E}(\mathcal{P})$, an object detector $\mathcal{D}$, training dataset $S$ and ground truth box $y$.
Parameter: DDIM steps $T$, unconditional embeddings iteration number $N_{u}$, training batch $B_{t}$, and patch scale $\tau$.
Output: Adversarial patch
1:  Compute $z^{*}_{\frac{T}{2}},\ldots,z^{*}_{0}$ using DDIM inversion over $\mathcal{I}^{\prime}$ with guidance scale $w=1$
2:  Set $w=7.5$, $\bar{z}_{\frac{T}{2}}\leftarrow z^{*}_{\frac{T}{2}}$, $\phi_{\frac{T}{2}}\leftarrow\mathcal{E}(""),\delta_{0}\leftarrow 0$
3:  for $t=\frac{T}{2},\frac{T}{2}-1,\ldots,1$ do
4:     for $j=0,\ldots,N_{u}-1$ do
5:        $\phi_{t}\leftarrow\phi_{t}-\eta_{u}\nabla_{\phi}\|{z^{*}_{t-1}-z_{t-1}(\bar{z_{t}},\phi_{t},\mathcal{C})}\|_{2}^{2}$
6:     end for
7:     $\bar{z}_{t-1}\leftarrow z_{t-1}(\bar{z_{t}},\phi_{t},\mathcal{C})$, $\phi_{t-1}\leftarrow\phi_{t}$
8:  end for
9:  $m^{*}\leftarrow Downsample(m)$
10:  for $k=1,\ldots,N_{iter}$ do
11:     for $b=1,\ldots,B_{t}$ do
12:        $X\leftarrow$ Extract the $b$-th batch of data from $S$
13:        $z_{p}\leftarrow G(\bar{z}_{\frac{T}{2}}+\delta_{b-1},\frac{T}{2},\mathcal{C},\{\phi_{t}\}^{\frac{T}{2}}_{t=1})\odot m$
14:        $X^{\prime}\leftarrow\mathcal{T}(z_{p},X,\tau)$
15:        $g_{b}\leftarrow Adam(\nabla_{\bar{z}_{\frac{T}{2}}}\mathcal{L}_{IoU}(\mathcal{D}(X^{\prime}),y))$
16:        $\delta_{b}\leftarrow Proj_{\infty}(\delta_{b-1}+\eta_{p}\cdot g_{b},0,\epsilon)\odot m^{*}$
17:     end for
18:     Set $\delta_{0}=\delta_{B}$
19:  end for
20:  $z_{p}\leftarrow G(\bar{z}_{\frac{T}{2}}+\delta_{B},\frac{T}{2},\mathcal{C},\{\phi_{t}\}^{\frac{T}{2}}_{t=1})\odot m$
21:  return $z_{p}$
4 Experiments
[…]
Method YLv3 YLv3t YLv4 YLv4t YLv5s YLv7t FRCNN DETR
AP |||| ASR AP |||| ASR AP |||| ASR AP |||| ASR AP |||| ASR AP |||| ASR AP |||| ASR AP |||| ASR
${}^{(P_{A})}$dvYL 31.9 |||| 46.3 13.6 |||| 66.4 46.3 |||| 36.2 9.0 |||| 69.0 24.8 |||| 53.4 61.8 |||| 20.7 38.3 |||| 24.0 33.1 |||| 36.7
${}^{(P_{B})}$-SEA 22.0 |||| 39.0 7.9 |||| 59.0 18.4 |||| 48.7 12.2 |||| 56.8 10.2 |||| 49.3 46.3 |||| 26.6 20.0 |||| 27.1 21.8 |||| 36.3
${}^{(P_{C})}$ 50.4 |||| 33.7 15.9 |||| 74.2 61.3 |||| 26.2 13.7 |||| 78.9 41.0 |||| 41.4 69.0 |||| 17.5 55.5 |||| 25.1 43.3 |||| 41.5
${}^{(P_{D})}$2D 50.3 |||| 33.9 18.8 |||| 69.1 58.4 |||| 30.9 15.6 |||| 75.0 41.0 |||| 40.7 68.7 |||| 18.4 52.9 |||| 25.0 44.8 |||| 34.2
${}^{(P_{1})}$3(Ours) 12.9 |||| 70.4 17.0 |||| 64.1 58.2 |||| 28.1 21.3 |||| 60.4 36.0 |||| 52.2 60.2 |||| 24.6 38.6 |||| 29.3 38.6 |||| 38.6
${}^{(P_{2})}$3t(Ours) 40.2 |||| 38.6 3.9 |||| 86.1 62.5 |||| 24.4 16.9 |||| 65.3 40.6 |||| 44.8 59.4 |||| 24.1 44.1 |||| 29.5 40.4 |||| 39.6
${}^{(P_{3})}$4(Ours) 32.3 |||| 45.0 23.8 |||| 57.5 35.1 |||| 49.9 31.1 |||| 46.8 45.1 |||| 43.5 62.4 |||| 21.6 43.7 |||| 27.4 39.0 |||| 39.0
${}^{(P_{4})}$4t(Ours) 40.8 |||| 41.0 11.8 |||| 73.6 61.6 |||| 25.5 9.2 |||| 79.1 38.3 |||| 50.3 57.4 |||| 25.7 45.3 |||| 28.4 39.5 |||| 41.4
${}^{(P_{5})}$5s(Ours) 33.9 |||| 47.4 17.8 |||| 64.3 60.3 |||| 27.2 23.7 |||| 56.5 22.0 |||| 69.9 61.3 |||| 23.1 40.6 |||| 29.9 37.2 |||| 38.3
${}^{(P_{6})}$7t(Ours) 49.6 |||| 35.0 19.9 |||| 64.3 64.2 |||| 23.0 26.0 |||| 56.7 39.6 |||| 44.5 51.7 |||| 33.5 50.2 |||| 26.1 42.2 |||| 38.9
${}^{(P_{7})}$(Ours) 38.8 |||| 42.7 19.9 |||| 62.4 58.1 |||| 29.2 23.0 |||| 60.8 37.2 |||| 50.7 61.7 |||| 24.3 22.6 |||| 38.5 35.4 |||| 43.7
${}^{(P_{8})}$(Ours) 46.6 |||| 35.2 35.2 |||| 49.0 67.1 |||| 22.1 34.3 |||| 46.7 44.6 |||| 43.1 66.8 |||| 20.5 46.2 |||| 26.1 37.1 |||| 42.0
${}^{(P_{j})}$ource 53.5 |||| 29.7 45.9 |||| 38.1 70.0 |||| 18.5 47.8 |||| 36.9 55.8 |||| 35.0 69.4 |||| 18.4 53.5 |||| 23.3 47.3 |||| 33.6
       Gray 72.3 |||| 16.5 61.8 |||| 25.2 76.8 |||| 12.8 56.4 |||| 30.0 74.1 |||| 15.5 77.1 |||| 13.5 64.4 |||| 17.3 50.2 |||| 34.4
${}^{(P_{m})}$andom 70.2 |||| 16.9 64.3 |||| 22.2 78.5 |||| 12.2 60.4 |||| 26.9 75.1 |||| 15.5 79.8 |||| 12.3 67.3 |||| 17.6 51.1 |||| 31.7
${}^{(P_{n})}$andom 72.8 |||| 17.0 68.0 |||| 19.0 82.4 |||| 10.4 69.8 |||| 20.7 83.2 |||| 11.7 79.3 |||| 12.4 72.0 |||| 16.4 57.0 |||| 25.9
$P_{1}$ $P_{2}$ $P_{3}$ $P_{4}$ $P_{5}$ $P_{6}$ $P_{7}$ $P_{8}$ $P_{A}$ $P_{B}$ $P_{C}$ $P_{D}$ $P_{j}$ $P_{m}$ $P_{n}$
4.1 Experimental Setup
Datasets
In our experiments, we trained and evaluated the adversarial patches using the INRIA person dataset [11], which consists of 614 training images and 288 test images. To meet the input requirements of the detector, we padded all images with gray pixels to make them square and then resized them to 640 $\times$ 640. Direct resizing would distort the image proportions. Furthermore, we evaluated the transferability of different adversarial patch methods on the MPII Human Pose dataset [1], COCO dataset [31] and 1,000 images from diverse scenarios collected from the Internet.
Attacked Models
Our experiments attack a variety of detectors, including YOLOv3, YOLOv3-tiny [37], YOLOv4, YOLOv4-tiny [4], YOLOv5s [25], YOLOv7-tiny [44], FasterRCNN [38] and DETR [6]. We will abbreviate YOLO as YL and FasterRCNN as FRCNN for ease of presentation, such as YOLOv4-tiny as YLv4t. All models are pre-trained on COCO dataset [31].
Implementation Details
We treat person as the target attack class, set the batch size to 32, the maximum epoch $N_{iter}=200$, DDIM sampling steps $T=50$, unconditional embedding optimization iterations $N_{u}=10$, $\eta_{u}=0.01$, $\eta_{p}=0.003$, $\epsilon=0.5$, we use Adam optimizer [26] with the default parameters, the textual prompts $\mathcal{P}$ are automatically generated using BLIP [29], the version of Stable Diffusion [39] is v1.4. Since we are generating irregular patches, we set the patch size to match the pixel area of a square patch with a scale of $\tau=0.2$. To achieve stronger performance in patch attacks, we applied two iterations of optimization to the patch using DiffPatch. We regard the prediction of detector on clean data as the ground truth, and measure the capability of adversarial patch attack using the Average Precision (AP). A lower AP indicates that the accuracy of the detector in recognizing persons is reduced, indirectly indicating stronger patch attack performance. Additionally, since AP only measures the overall decline trend caused by the patches and does not directly reflect the attack performance, we also evaluated the Attack Success Rate (ASR) of various patches. The confidence threshold and IoU threshold for the detector are both set to 0.5. All of our experiments were conducted on a single NVIDIA Tesla A100.
4.2 Attack Performance Evaluation
We generated a series of adversarial patches targeting different models using DiffPatch and compared them with current patch generation methods on INRIA dataset, including AdvYL [43], T-SEA [22], NPAP [21], and D2D [30]. The detailed results can be seen in Table 1. DiffPatch ($P_{1}-P_{8}$) performs comparably to the best existing non-natural patch generation methods in terms of the AP metric. Notably, DiffPatch achieved the best Attack Success Rate (ASR), with an improvement of 10.1% over AdvYL on YLv4t. Compared to naturalistic adversarial patches ($P_{C}$ and $P_{D}$), our patch trained on YLv4t effectively reduced the AP by 4.5% and 6.4%, and increased the ASR by 0.2% and 4.1%.
4.3 Cross-dataset Evaluation
Considering the relatively simple scenes in the INRIA dataset, we conducted tests to evaluate the attack performance of the patches in more diverse environments. We trained patches using YLv4t with various methods and tested them on multiple datasets, including MPII and COCO. Due to the large size of these datasets, we randomly sampled a subset of images containing people for testing, selecting 729 images from COCO and 569 images from MPII. Additionally, since the images from these two datasets are relatively outdated with low-resolution images, we also collected 1,000 high-resolution images from the Internet, representing diverse scenes (denoted as “Collected”). The results, as shown in Table 2, demonstrate that our method significantly reduces the AP while substantially increasing the ASR. On average, the ASR shows a 17.5% improvement over the unnatural patch method T-SEA and a 2.6% improvement over the natural patch method NPAP.
Metric Dataset Attack
T-SEA NPAP D2D Ours
AP%$\downarrow$ INRIA 12.2 13.7 15.6 9.2
MPII 17.0 17.3 17.9 12.5
COCO 21.4 25.1 26.8 16.9
Collected 8 6.7 7.2 3.9
ASR%$\uparrow$ INRIA 56.8 78.9 75.0 79.1
MPII 60.3 73.4 73.3 75.9
COCO 54.5 63.9 60.3 67.8
Collected 67.2 82.1 79.4 85.8
Table 2: Attack performance of different attacks across various datasets.
4.4 Ensemble Attack Evaluation
We employed an ensemble approach, training our patch using the averaged gradients from multiple models. Compared to other natural patch generation methods, our single patch achieved strong attack performance across various object detection models. As shown in Table 3, NPAP was trained on YLv3 and YLv4t, while D2D and our DiffPatch were trained on all models. In white-box attacks, our method demonstrated a significant performance improvement over D2D.
[…]
Method YLv3 YLv3t YLv4 YLv4t YLv5s YLv7t FRCNN Avg.
AP |||| ASR AP |||| ASR AP |||| ASR AP |||| ASR AP |||| ASR AP |||| ASR AP |||| ASR AP |||| ASR
NPAP 45.2 |||| 34.2 16.2 |||| 74.8 65.8 |||| 22.6 19.3 |||| 71.7 57.4 |||| 28.5 70.8 |||| 17.3 57.8 |||| 24.0 32.3 |||| 53.0
D2D 48.9 |||| 35.7 17.5 |||| 66.8 58.4 |||| 28.9 24.6 |||| 59.3 35.6 |||| 43.2 67.0 |||| 19.5 52.7 |||| 24.3 43.5 |||| 39.7
DiffPatch 9.1 |||| 65.9 6.1 |||| 79.5 28.9 |||| 47.4 12.0 |||| 75.5 23.2 |||| 62.4 55.4 |||| 28.5 19.9 |||| 38.9 22.1 |||| 56.9
4.5 Ablation Studies
Loss Function
We conducted experiments using both the Common Detection Loss and IoU-Detection Loss under the same patch training setting. As shown in Figure 4, compared to the Common Detection Loss, our IoU-Detection Loss demonstrates greater stability during training process and achieves higher attack performance in a earlier time.
Figure 4: Comparison of two loss functions. Left: training loss; Right: attack performance.
Shape Impact
The impact of patch shapes was investigated, as indicated by the data for $P_{m}$ and $P_{n}$ in Table 1. We compared the performance of random noise images with square and irregular shapes (with the same size as DiffPatch). From the AP and ASR metrics, it is evident that square patches yield higher attack benefits compared to irregular patches. In other words, designing irregular patches with high attack success rates is more challenging.
Diffusion Trajectory Length
Using latent vectors from intermediate timestamps to generate patches effectively mitigates semantic loss. We tested the results of latent vectors at different timestamps, evaluating the attack performance and similarity to textual descriptions for patches generated from diffusion trajectories of varying lengths. See Figure 5 for details. We use the cosine similarity of CLIP [35] embeddings to measure the alignment between the patch and the prompt, which indirectly reflects the visual quality of the patch. It is evident that starting diffusion optimization from $\frac{T}{2}$ (25 Steps) not only achieves a high success rate but also produces images that better align with textual descriptions, thereby maintaining greater semantic consistency with the original images.
Figure 5: The effects of diffusion trajectories with different lengths on the performance of patch attacks and natural appearance.
Target Mask
The target mask mechanism helps mitigate semantic loss in images to some extent. We compared the training results with and without the use of this mechanism. Without using target mask leads to quality issues in the patches during optimization, as shown in Figure 6.
Figure 6: The impact of target mask on adversarial patch optimization. Without target mask control, adversarial patches exhibit noticeable detail distortions (indicated by red boxes).
Optimization Constraint
We use the infinity norm to constrain each optimization of the patch to ensure its natural appearance. The size of $\epsilon$ significantly impacts patch generation. An appropriate $\epsilon$ can effectively balance the trade-off between attack performance and appearance. As shown in Table 4, we ultimately chose $\epsilon=0.5$ as the constraint parameter for patch generation, as it maintains both appearance and a high attack success rate.
Constraint $\epsilon$ 0.3 0.5 0.7
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
CLIP Sim. 75.6% 73.2% 63.9%
ASR 57.0% 63.0% 67.4%
Table 4: The effect of optimization constraint $\epsilon$ on latent vectors.
Iterative Optimization
As the results in Table 5, iterative optimization can further enhance the adversarial strength of the patch. However, it also makes the patch unnatural.
Metric Iterative Model
YLv3t YLv4t YLv5s FRCNN
AP%$\downarrow$ w/o 7.1 12.3 24.6 26.0
w/ 3.9 9.2 22.0 22.6
ASR%$\uparrow$ w/o 82.9 75.5 65.6 37.3
w/ 86.1 79.1 69.9 38.5
Sim.$\uparrow$ w/o 77.1 77.6 73.2 72.9
w/ 72.1 69.4 60.0 60.8
Table 5: The impact of using iterative optimization.
4.6 Physical-world Adversarial Patch Dataset
Our DiffPatch achieves promising attack results even in physical-world environments. Using DiffPatch, we designed 9 adversarial patches and printed them on T-shirts, as shown in Figure 7. We recruited 20 participants (with ethics approval) to capture images in diverse indoor and outdoor, including laboratories, campus, cafeteria, subway station, and shopping mall. In total, we collected 1,131 images into our AdvPatch-1K dataset, including both individual and group (2 to 10+ persons) photos. The dataset is accompanied with detailed annotations of the person and patch locations (bounding boxes). Several example images are shown in Figure 1, with detailed statistics in Figure 8. The images in AdvPatch-1K can effectively evade the YOLOv5s detector, as shown in the right subfigure of Figure 8. Our dataset provides a valuable resource for advancing defense research against adversarial patch attacks.
Figure 7: Adversarial T-shirts created using DiffPatch.
Figure 8: Left: The proportion of images with different numbers of persons in AdvPatch-1K. Right: The detection performance of YOLOv5s on AdvPatch-1K under different confidence thresholds.
5 Conclusion
In this work, we proposed a novel diffusion-based adversarial patch generation method DiffPatch to generate naturalistic and customized patches based on a reference image. DiffPatch utilizes image-to-latent mapping, incomplete diffusion optimization, and target mask control to create stylized adversarial patches. Extensive experiments across multiple object detection models validate the effectiveness of DiffPatch. With DiffPatch, we also created a physical-world adversarial dataset, AdvPatch-1K, comprising 1,131 images captured in diverse scenes and conditions. AdvPatch-1K can help develop robust defense methods against adversarial patch attacks.

### v3

@@
-DiffPatch: Generating Customizable Adversarial Patches using Diffusion Model
+DiffPatch: Generating Customizable Adversarial Patches using Diffusion Models
 Abstract
-Physical adversarial patches printed on clothing can easily allow individuals to evade person detectors.
-However, most existing adversarial patch generation methods prioritize attack effectiveness over stealthiness, resulting in patches that are aesthetically unpleasing.
-Although existing methods using generative adversarial networks or diffusion models can produce more natural-looking patches, they often struggle to balance stealthiness with attack effectiveness and lack flexibility for user customization.
-To address these challenges, we propose a novel diffusion-based customizable patch generation framework termed DiffPatch, specifically tailored for creating naturalistic and customizable adversarial patches.
-Our approach enables users to utilize a reference image as the source, rather than starting from random noise, and incorporates masks to craft naturalistic patches of various shapes, not limited to squares.
-To prevent the original semantics from being lost during the diffusion process, we employ Null-text inversion to map random noise samples to a single input image and generate patches through Incomplete Diffusion Optimization (IDO).
-Notably, while maintaining a natural appearance, our method achieves a comparable attack performance to state-of-the-art non-naturalistic patches when using similarly sized attacks.
-Using DiffPatch, we have created a physical adversarial T-shirt dataset, AdvPatch-1K, specifically targeting YOLOv5s.
-This dataset includes over a thousand images across diverse scenarios, validating the effectiveness of our attack in real-world environments.
-Moreover, it provides a valuable resource for future research.
+Physical adversarial patches printed on clothing can enable individuals to evade person detectors, but most existing methods prioritize attack effectiveness over stealthiness, resulting in aesthetically unpleasing patches.
+While generative adversarial networks and diffusion models can produce more natural-looking patches, they often fail to balance stealthiness with attack effectiveness and lack flexibility for user customization.
+To address these limitations, we propose DiffPatch, a novel diffusion-based framework for generating customizable and naturalistic adversarial patches.
+Our approach allows users to start from a reference image (rather than random noise) and incorporates masks to create patches of various shapes, not limited to squares.
+To preserve the original semantics during the diffusion process, we employ Null-text inversion to map random noise samples to a single input image and generate patches through Incomplete Diffusion Optimization (IDO).
+Our method achieves attack performance comparable to state-of-the-art non-naturalistic patches while maintaining a natural appearance.
+Using DiffPatch, we construct AdvT-shirt-1K, the first physical adversarial T-shirt dataset comprising over a thousand images captured in diverse scenarios.
+AdvT-shirt-1K can serve as a useful dataset for training or testing future defense methods.
 1 Introduction
-Figure 1: Example images from our AdvPatch-1K dataset, showcasing the performance of DiffPatch in various scenarios.
-With the rapid advancement of computer vision, object detectors are extensively utilized in real-world scenarios, including fields such as autonomous driving [8] and medical diagnostics [33].
-Unfortunately, recent works [7, 47, 15, 45] have demonstrated that deep neural networks are vulnerable to maliciously crafted adversarial examples, resulting in decisions that defy common sense.
-This poses significant threats to safety-critical applications.
-In the field of computer vision, adversarial attacks can be categorized into two distinct groups: digital attacks [17, 7] and physical attacks [14, 27, 32].
-Different from digital attacks, which involve introducing small perturbations within a digital domain, physical adversarial attacks operate in the tangible world, aiming to mislead visual systems through modifications to the appearance or positioning of physical objects.
-The execution of physical adversarial attacks can be accomplished through numerous methods, including the application of adversarial patches onto the surface of an object, alterations in environmental lighting conditions, or the introduction of distracting elements.
-For physical-world adversarial patches, previous research [43, 48] has primarily focused on enhancing attack performance, which often requires introducing stronger perturbations that significantly deviate patches from the distribution of natural images, making them appear highly unnatural.
-While most patch generation methods [46, 23] incorporate a total variation loss to smooth the patches, this fundamental issue remains unresolved, rendering the patches easily identifiable by humans.
-With the rapid development of image generation techniques [16, 20], naturalistic patch generation methods based on deep image generators have gradually emerged [21, 30].
-The patches generated by sate-of-the-art generators possess a more natural appearance, significantly enhancing their stealthiness.
-However, these methods suffer from limited attack effectiveness compared to unnatural patches.
-Moreover, these methods have high randomness in the generation process, making it difficult to generate customized patches based on a given image.
-In this work, we propose a novel Stable Diffusion [39] based adversarial patch generation method called DiffPatch that is capable of generating a naturalistic and customizable adversarial patch based on a reference image specified by the adversary.
-Our DiffPatch method works as follows.
-Firstly, it maps the reference image into the latent space and generates adversarial patches by optimizing the noise latent vectors.
-It adopts Null-text inversion [34] to achieve a near-perfect reconstruction so as to guarantee that the generated images largely preserve the original semantics.
-Secondly, it introduces an Incomplete Diffusion Optimization (IDO) strategy to ensure the natural appearance and attack performance of the optimized patches and employs an IoU-Detection loss to accelerate convergence.
-Furthermore, to generate irregular-shaped patches, it utilizes masks to replace the background and covers the gradients in non-essential areas.
-This can also help reduce information interference from other regions during the optimization process.
-Based on DiffPatch, we have created an adversarial patch dataset AdvPatch-1K which consists of over a thousand images, featuring attacks in various physical-world scenarios with customizable patches of diverse styles and shapes.
-Partial results are shown in Figure 1.
-Our main contributions are as follows:
+Figure 1: Example images from our AdvT-shirt dataset, showcasing the performance of DiffPatch in various scenarios.
+Object detectors powered by deep neural networks (DNNs) are integral to modern computer vision systems, enabling critical applications such as autonomous driving [9] and medical diagnostics [39].
+However, DNNs are highly susceptible to adversarial attacks [8, 57, 16, 55], which manipulate their outputs in ways that defy logic or compromise safety.
+These attacks can cause object detectors to misclassify or fail to detect critical objects.
+This vulnerability poses significant risks, particularly in safety-critical domains.
+Adversarial attacks on vision models can be broadly categorized into digital attacks [18, 8] and physical attacks [15, 30, 37, 32, 54].
+Unlike digital attacks, which introduce imperceptible perturbations in the digital domain, physical attacks manipulate real-world objects or environments to deceive vision models.
+Common physical attack strategies include applying adversarial patches [50], altering lighting conditions [62], or introducing distracting elements [36].
+While adversarial patches have shown promise in physical attacks, most existing methods [50, 58] prioritize attack effectiveness at the expense of stealthiness, resulting in patches that appear unnatural and are easily detectable.
+Although techniques like total variation loss [56, 26] have been employed to smooth patches, they fail to fully address this issue.
+Recent advances in image generation [17, 23] have enabled the creation of more naturalistic patches [24, 34], but these methods often sacrifice attack effectiveness and lack customization capabilities.
+For instance, recent work [11] leverages image-latent mapping for adversarial sample generation, but its large perturbations in physical settings lead to poor naturalness or effectiveness.
+To address these limitations, we propose DiffPatch, a novel adversarial patch generation framework based on Stable Diffusion [45].
+DiffPatch generates naturalistic and customizable adversarial patches from a reference image specified by the adversary.
+Our approach first maps the reference image into the latent space and optimizes noise latent vectors to create adversarial patches.
+We employ Null-text inversion [40] to ensure near-perfect reconstruction, preserving the original semantics.
+Additionally, we introduce an Incomplete Diffusion Optimization (IDO) strategy to balance natural appearance and attack performance, along with an IoU-Detection loss to accelerate convergence.
+To generate irregular-shaped patches, we use masks to replace the background and suppress gradients in non-essential areas, reducing interference during optimization.
+Using DiffPatch, we construct AdvT-shirt-1K, a dataset of over a thousand images featuring customizable adversarial patches in diverse physical-world scenarios.
+This dataset demonstrates the practicality of our approach and provides a valuable resource for future research.
+In summary, our key contributions are as follows:
   • We propose a novel diffusion-based method called DiffPatch for generating naturalistic adversarial patches based on a reference image.
-DiffPatch employs Null-text inversion, Incomplete Diffusion Optimization (IDO) , masked customization, and an IoU-Detection loss to generate stylized and customizable adversarial patches.
-  • Experimental results show that our DiffPatch achieves the highest attack success rate (ASR), significantly reducing the mAP of various object detectors, rivaling or even surpassing the current state-of-the-art methods of unnatural adversarial patches.
-  • We create physical-world adversarial patches by printing them on T-shirts and collect a dataset, AdvPatch-1K, which consists of 1,131 images with annotations.
-This dataset covers 9 unique adversarial T-shirt designs across various contexts, including both indoor and outdoor settings as well as individual and group photos.
+DiffPatch integrates Null-text inversion, Incomplete Diffusion Optimization (IDO) , masked customization, and an IoU-Detection loss to generate stylized and customizable adversarial patches.
+  • Experimental results demonstrate that our DiffPatch achieves the highest attack success rate (ASR), significantly reducing the mean average precision (mAP) of various object detectors.
+Its performance rivals or even surpasses state-of-the-art methods for generating unnatural adversarial patches.
+  • We construct AdvT-shirt-1K, a dataset of physical-world adversarial patches printed on T-shirts.
+It includes 1,131 annotated images featuring 9 unique adversarial designs across diverse scenarios, including indoor and outdoor environments, as well as individual and group photos, providing a valuable resource for future defense research.
 2 Related Work
-2.1 Physical Adversarial Attack
-With the widespread application of deep learning, researchers began to investigate the security of these models in real-world scenarios.
-This focus led to the development of the Basic Iterative Method (BIM) [27], which generates adversarial samples that retain their adversarial properties even after being printed.
-To improve the robustness of physical adversarial samples, the Expectation Over Transformation (EOT) approach [2] was introduced.
-EOT enhances resistance to real-world transformations by simulating various modifications that images may undergo.
-The Digital-to-Physical (D2P) method [24] further advances real-world attack effectiveness by learning the transformations between digital and physical images.
-AdvPatch [5] marks a milestone by demonstrating the creation of robust adversarial patches that function reliably across diverse environmental conditions.
-Following this, AdvYOLO [43] demonstrated an attack on object detectors by applying adversarial patches to clothing surfaces, allowing wearers to evade detection.
-Several similar patch attack methods have since been proposed [48, 46, 52, 22], posing a significant threat to object detectors.
-Additionally, other physical adversarial attack methods have employed accessories like glasses [41], lighting [51], and shadows [50].
+2.1 Physical Adversarial Attacks
+Early research on adversarial attacks primarily focused on digital domains, with methods like Basic Iterative Method (BIM) [30] and Projected Gradient Descent (PGD) [38] generating perturbations that remain effective even after physical transformations.
+To address real-world challenges, the Expectation Over Transformation (EOT) approach [3] was introduced, simulating environmental variations to enhance robustness.
+The Digital-to-Physical (D2P) method [27] further improved attack effectiveness by modeling the transition between digital and physical domains.
+These advancements laid the groundwork for physical adversarial attacks, such as AdvPatch [6], which demonstrated robust adversarial patches effective across diverse conditions.
+Subsequent work, including AdvYOLO [50], extended these ideas to evade object detectors using clothing-based patches.
+Other physical attack strategies have leveraged accessories like glasses [47], manipulated lighting [62], and exploited shadows [61], highlighting the expanding scope of real-world adversarial threats.
 2.2 Naturalistic Adversarial Patch
-Inspired by the imperceptibility constraints in digital attacks, Universal Physical Camouflage Attacks (UPC) [23] employ the $L_{\infty}$ norm to limit perturbations and maintain the natural appearance of patches.
-Unlike iterative methods that optimize patches on images, the Naturalistic Adversarial Patch (NAP) [21] utilizes a generative approach.
-NAP leverages GANs to indirectly optimize the patch by modifying initial latent vectors, resulting in patches with a highly natural appearance.
-Building on these advancements, several GAN-based generation methods have emerged [13, 28].
-Additionally, diffusion model-based methods have gained traction in the vision field [9, 30, 49].
-The generative power of diffusion models enables the creation of more naturalistic and effective physical adversarial patches.
+Drawing inspiration from the imperceptibility constraints in digital attacks, Universal Physical Camouflage Attacks (UPC) [26] use the $L_{\infty}$ norm to bound perturbations, ensuring patches retain a natural appearance.
+Unlike traditional iterative methods that optimize patches directly on images, the Naturalistic Adversarial Patch (NAP) [24] adopts a generative approach.
+NAP employs GANs to optimize patches indirectly by modifying initial latent vectors, producing patches that are both natural-looking and effective.
+This approach has spurred the development of several GAN-based methods [14, 31].
+More recently, diffusion models have emerged as a powerful tool in the vision domain [10, 34, 59], leveraging their generative capabilities to create even more naturalistic and effective physical adversarial patches.
+These advancements highlight the growing potential of generative models in crafting stealthy and robust adversarial attacks.
 Figure 2: An overview of the proposed DiffPatch generation framework.
 2.3 Diffusion-based Image Editing
-Recently, diffusion models have excelled in the field of image generation, with models like DALL-E2 [36], Imagen [40], and Stable Diffusion [39] largely surpassing GANs in performance [12].
-Beyond generation tasks, diffusion-based image editing techniques have also become mainstream.
-Leveraging the cross-attention mechanism, Prompt to Prompt (P2P) [18] introduces an innovative framework that enables image editing solely through text prompts.
-To address challenges with poor reconstruction during editing, Null-text inversion [34] achieves near-perfect reconstruction by optimizing the unconditional embedding at each denoising step.
-Text-driven Blended Diffusion [3] employs masks to target specific regions in an image, allowing for background replacement and consistent edits within defined areas.
-Additionally, some adversarial attack methods [49, 10] leverage image editing techniques to modify images while preserving their natural appearance, creating adversarial examples that are visually coherent yet misleading to detection models.
+Diffusion models have recently achieved remarkable success in image generation, with models like DALL-E2 [42], Imagen [46], and Stable Diffusion [45] outperforming GANs in various benchmarks [13].
+Beyond generation, diffusion-based image editing techniques have gained prominence [48].
+For instance, Prompt to Prompt (P2P) [21] leverages the cross-attention mechanism to enable text-driven image editing, allowing precise modifications through textual prompts.
+To address reconstruction challenges during editing, Null-text inversion [40] achieves near-perfect reconstruction by optimizing unconditional embeddings at each denoising step.
+Text-driven Blended Diffusion [4] uses masks to target specific regions, enabling background replacement and localized edits while maintaining consistency.
+These advancements have also inspired adversarial attack methods [59, 11], which utilize image editing techniques to craft visually coherent adversarial examples that deceive detection models while preserving natural appearance.
 3 Proposed Approach
+In this section, we begin with an overview of our proposed attack framework, followed by a detailed introduction to its three key components.
 Overview
-Figure 2 provides an overview of DiffPatch.
-It optimizes the Null text embeddings along incomplete diffusion trajectories and maps the image to the latent vectors at intermediate timestamps, ensuring that the patch maintains a natural appearance during updates.
-It then applies the Incomplete Diffusion Optimization (IDO) guided by an IoU-Detection Loss to optimize the latent vectors, resulting in faster convergence and enhanced attack performance.
-During optimization, we skip the U-Net gradients to avoid memory overflow and utilize target masks to eliminate semantic interference from irrelevant areas, thereby generating more naturalistic adversarial patches.
-Next, we will introduce the key techniques used in DiffPatch.
+Figure 2 illustrates the overall framework of DiffPatch.
+It optimizes Null-text embeddings along incomplete diffusion trajectories and maps the image to latent vectors at intermediate timestamps, ensuring that the patch retains a natural appearance during updates.
+It employs Incomplete Diffusion Optimization (IDO) guided by an IoU-Detection Loss to optimize the latent vectors, achieving faster convergence and improved attack performance.
+To address memory constraints, we skip the U-Net gradients during optimization and use target masks to eliminate semantic interference from irrelevant regions, enabling the generation of more naturalistic adversarial patches.
+Next, we will introduce the key techniques of DiffPatch.
 3.1 Image-to-Latent Mapping
-Our objective is to achieve a perfect mapping between a given image $\mathcal{I}$ and latent vector $z_{t}$, utilizing the image encoding $z_{0}$ to obtain $z_{t}$, which is essentially the reverse process of DDIM sampling [42].
+Our goal is to achieve a precise mapping between a given image $\mathcal{I}$ and its corresponding latent vector $z_{t}$.
+This is accomplished by leveraging the image encoding $z_{0}$ to obtain $z_{t}$, which is essentially the reverse process of DDIM sampling [49].
 Based on the assumption that the ODE process can be reversed in the limit of small steps, DDIM inversion can be formulated as:
 $z_{t+1}=\sqrt{\frac{\alpha_{t+1}}{\alpha_{t}}}z_{t}+\left(\sqrt{\frac{1}{\alpha_{t+1}}-1}-\sqrt{\frac{1}{\alpha_{t}}-1}\right)\cdot\varepsilon_{\theta}(z_{t},t,\mathcal{C}),$
-where $\varepsilon_{\theta}$ is the pre-trained U-Net used to predict noise, $\mathcal{C}$ is the condition embedding of $\mathcal{P}$, and $\alpha_{t}$ is calculated based on the schedule $\beta_{0},\ldots,\beta_{T}\in(0,1)$, i.e., $\alpha_{t}=\prod_{i=1}^{t}(1-\beta_{i})$.
-Through the aforementioned formula, we can obtain the latent vector $z_{t}$ and utilize it for image reconstruction, yet this does not guarantee perfect reconstruction.
-To enhance the quality of generated outputs, Stable Diffusion incorporates a classifier-free guidance mechanism [19] that concurrently engages in unconditional generation and guided generation based on condition embedding $\mathcal{C}$:
+where $\varepsilon_{\theta}$ is the pre-trained U-Net for noise prediction, $\mathcal{C}$ is the condition embedding of $\mathcal{P}$, and $\alpha_{t}$ is calculated based on the schedule $\beta_{0},\ldots,\beta_{T}\in(0,1)$, i.e., $\alpha_{t}=\prod_{i=1}^{t}(1-\beta_{i})$.
+While this formula allows us to obtain the latent vector $z_{t}$ and reconstruct the image, it does not guarantee perfect reconstruction.
+To improve the quality of generated outputs, Stable Diffusion employs a classifier-free guidance mechanism [22] , which combines unconditional generation with guided generation based on condition embedding $\mathcal{C}$:
 $\tilde{\varepsilon}_{\theta}(z_{t},t,\mathcal{C},\phi)=w\cdot\varepsilon_{\theta}(z_{t},t,\mathcal{C})+(1-w)\ \cdot\varepsilon_{\theta}(z_{t},t,\phi),$ (2)
-where $w$ represents the guidance scale parameter, and $\phi=\mathcal{E}("")$ is the embedding of a null text.
-The generated content is highly sensitive to textual prompts.
-During each step of guided generation, even minor variations in the prompt can lead to deviations in the trajectory.
-This cumulative error can make reconstruction challenging, and such errors tend to amplify when we perform image editing.
-To resolve this issue, we separately optimize the null text embedding $\phi_{t}$ for each timestamp $t$ in the order of the diffusion process $t=T\rightarrow t=1$, using the optimized embedding to guide the process back onto the correct trajectory.
-Similar to Null-text inversion [34], firstly, we set the guidance scale $w$ to 1 and utilize DDIM inversion to obtain intermediate vectors $z^{*}_{T},\ldots,z^{*}_{0}$ where $z^{*}_{0}=z_{0}$.
-After that, regarding $z^{*}_{T}$ as the initial noise vector $\bar{z}_{T}$, we optimize the null text embedding $\phi_{t}$ through the application of the following objective function, after each iteration of the DDIM sampling step (adopting the default value of 7.5 for the guidance scale $w$):
+where $w$ is the guidance scale parameter, and $\phi=\mathcal{E}("")$ is the embedding of a null text.
+The generated content is highly sensitive to textual prompts, and even minor variations can cause trajectory deviations during guided generation.
+These deviations accumulate, making reconstruction challenging, especially during image editing.
+To address this, we separately optimize the null text embedding $\phi_{t}$ for each timestamp $t$ in the diffusion process $t=T\rightarrow t=1$, guiding the process back to the correct trajectory.
+Similar to Null-text inversion [40], we first set the guidance scale $w$ to 1 and use DDIM inversion to obtain intermediate vectors $z^{*}_{T},\ldots,z^{*}_{0}$, where $z^{*}_{0}=z_{0}$.
+Treating $z^{*}_{T}$ as the initial noise vector $\bar{z}_{T}$, we optimize $\phi_{t}$ using the following objective function after each iteration of the DDIM sampling step (with the guidance scale $w$ set to the default value of 7.5):
 $\min_{\phi_{t}}\|z^{*}_{t-1}-z_{t-1}(\bar{z_{t}},\phi_{t},\mathcal{C})\|^{2}_{2},$ (3)
-where $z_{t-1}$ represents the vector obtained after DDIM sampling, and after each step of optimization is completed, we update $\bar{z}_{t-1}=z_{t-1}$.
-Finally, we can achieve near-perfect reconstruction by utilizing the initial noise vector $\bar{z}_{T}$ and the optimized unconditional embeddings $\{\phi_{t}\}^{T}_{t=1}$.
+where $z_{t-1}$ is the vector obtained after DDIM sampling.
+After each optimization step, we update $\bar{z}_{t-1}=z_{t-1}$.
+This process enables near-perfect reconstruction using the initial noise vector $\bar{z}_{T}$ and the optimized unconditional embeddings $\{\phi_{t}\}^{T}_{t=1}$.
 3.2 Incomplete Diffusion Optimization
-We propose Incomplete Diffusion Optimization (IDO), which adds adversarial perturbations along the trajectory of image reconstruction.
-IDO consists of two techniques: 1) IoU-Detection Loss and 2) Incomplete Diffusion Process.
+We propose Incomplete Diffusion Optimization (IDO), a method that introduces adversarial perturbations along the trajectory of image reconstruction.
+IDO comprises two key techniques: 1) IoU-Detection Loss and 2) Incomplete Diffusion Process.
 IoU-Detection Loss
-Loss function is crucial in guiding the optimization process.
-After an intermediate adversarial image is fed into the detector, we obtain the following outputs: 1) $B_{box}$: coordinates of the predicted bounding box; 2) $P_{obj}$: probability indicating the presence of an object within the predicted box; and 3) $P_{cls}$: classification probabilities for various objects.
-Previous work used Common Detection Loss [43] to guide adversarial patch generation:
+When an intermediate adversarial image is input into the detector, it produces three outputs: 1) $B_{box}$: the coordinates of the predicted bounding box; 2) $P_{obj}$: the probability of an object being present within the predicted box; and 3) $P_{cls}$: the classification probabilities for different objects.
+Previous work [50] employed the Common Detection Loss to guide adversarial patch generation:
 $\mathcal{L}_{det}=\frac{1}{N}\sum^{N}_{i=1}\max_{\begin{subarray}{c}j\end{subarray}}(P^{j}_{obj}(I^{\prime}_{i})\cdot P^{j}_{cls}(I^{\prime}_{i})),$ (4)
-where $N$ represents the batch size, and $I^{\prime}_{i}$ refers to the $i$-th image in a batch.
-For each image $I^{\prime}_{i}$, the detector identifies multiple objects and assigns an index $j$ to each detected object.
-The images $\{I^{\prime}_{i}\}_{i=1}^{N}$ are obtained by performing operation $\mathcal{T}$ on clean images, which means pasting patches on the surface of the clothes of the person inside.
-According to the loss function formulation in Equation (4), the optimization for a given image $I^{\prime}_{i}$ is guided solely by the object with the highest confidence product, aiming to minimize the detection probability for that specific target.
-However, this strategy introduces several challenges.
-First, the application of adversarial patches is constrained by the availability and accuracy of dataset labels.
-Incomplete annotations or misdetections can result in critical surfaces remaining unpatched, undermining the optimization process.
-Moreover, focusing the loss on a single object per iteration leads to slower convergence and increased instability during training, limiting the overall effectiveness of the approach.
-We propose an alternative loss function IoU-Detection Loss to address these issues. Formally, it is defined as:
+where $N$ is the batch size, and $I^{\prime}_{i}$ denote the $i$-th image in the batch.
+For each image $I^{\prime}_{i}$, the detector identifies multiple objects, assigning an index $j$ to each detection.
+The images $\{I^{\prime}_{i}\}_{i=1}^{N}$ are generated by applying an operation $\mathcal{T}$ on clean images, which involves pasting patches onto the clothing of individuals in the images.
+The loss function in Eq.(4) optimizes each image $I^{\prime}_{i}$ by focusing solely on the object with the highest confidence product, aiming to minimize its detection probability.
+However, this approach presents several limitations.
+First, the effectiveness of adversarial patches depends heavily on the availability and accuracy of dataset labels.
+Incomplete annotations or misdetections can leave critical surfaces unpatched, hindering the optimization process.
+Second, concentrating the loss on a single object per iteration slows convergence and increases training instability, reducing the overall effectiveness of the method.
+We propose an alternative loss function IoU-Detection Loss to address these limitations. Formally, it is defined as:
 $\mathcal{L}_{IoU}=\frac{1}{N}\sum_{i=1}^{N}\left(\frac{1}{M}\sum_{k=1}^{K}\left[\mathbbm{1}\left(\max_{j}\mathrm{IoU}(J_{j},J^{\prime}_{k})>t\right)P(J^{\prime}_{k})\right]\right),$
@@
 Incomplete Diffusion Process
-We found that using the default 50-step sampling to generate patches often leads to a loss of semantic meaning and natural appearance during optimization.
-Excessive sampling steps can amplify subtle changes in $z_{T}$, causing the image reconstruction trajectory to deviate.
-During the image-to-latent mapping phase, we can match it with the latent vector $z_{\frac{T}{2}}$ at an intermediate timestamp to mitigate semantic loss.
-The generation of adversarial patch $z_{p}$ can then be represented as:
+We observed that using the default 50-step sampling to generate patches often results in a loss of semantic meaning and a diminished natural appearance during optimization.
+Excessive sampling steps can amplify subtle changes in $z_{T}$, causing significant deviations in the image reconstruction trajectory.
+To address this, we employ an incomplete diffusion process, where the image is matched to the latent vector $z_{\frac{T}{2}}$ at an intermediate timestamp during the image-to-latent mapping phase.
+This approach reduces cumulative errors, preserving the patch’s semantic integrity and enhancing its natural appearance.
+The generation of the adversarial patch $z_{p}$ can then be represented as:
 $z_{p}=G(\bar{z}_{\frac{T}{2}},\frac{T}{2},\mathcal{C},\{\phi_{t}\}^{\frac{T}{2}}_{t=1}),$ (6)
 where $G$ represents the diffusion model.
-Generating patches using $z_{\frac{T}{2}}$ shortens the generation path, thereby enhancing the stability of the patches under adversarial guidance.
-However, as $z_{\frac{T}{2}}$ continues to be updated, it becomes more chaotic, making complete denoising challenging in the later stages of the process.
-To prevent adversarial perturbations from being noticeable, perturbation values are constrained by an infinity norm in digital attacks.
-Similarly, we constrain the perturbations $\delta$ applied to the latent vector $z_{\frac{T}{2}}$:
+By generating patches using $z_{\frac{T}{2}}$, we shorten the generation path, improving the stability of the patches under adversarial guidance.
+This method effectively minimizes semantic loss while maintaining naturalness.
+However, as $z_{\frac{T}{2}}$ is updated, it becomes more chaotic, making complete denoising increasingly challenging in later stages.
+To prevent noticeable adversarial perturbations, we constrain the perturbations $\delta$ applied to the latent vector $z_{\frac{T}{2}}$:
 $\delta_{t}=\text{Proj}_{\infty}(\delta_{t-1}+\Delta\delta,0,\epsilon),$ (7)
@@
 This ensures that the internal semantics of $z_{\frac{T}{2}}$ are not excessively lost, enabling the generation of patches with natural appearances through the denoising process.
-Specifically, during denoising, 25 steps of DDIM sampling are required, where U-Net performs 25 noise predictions.
-The denoising process involves removing noise at timestamp $t$ using a sampling recurrence formula to obtain the vector $z_{t-1}$, then gradually obtaining $z_{0}$, and finally using an image decoder $\Psi$ to restore the image.
-When optimizing the vector $z_{T/2}$ with the loss function $\mathcal{L}$, it can be expanded using the chain rule as follows:
-$\nabla_{z_{T/2}}\mathcal{L}=\frac{\partial\mathcal{L}}{\partial x}\cdot\frac{\partial x}{\partial z_{p}}\cdot\frac{\partial z_{p}}{\partial z_{0}}\cdot\frac{\partial z_{0}}{\partial z_{1}}\cdot\cdots\cdot\frac{\partial z_{T/2-1}}{\partial z_{T/2}},$ (8)
-where $z_{p}$ represents the adversarial patch obtained through the image encoder $\Psi$, i.e., $z_{p}=\Psi(z_{0})$ , and $x$ is the training data with $z_{p}$ attached.
-For each step of DDIM sampling, it can be expressed as:
+Specifically, during denoising, 25 steps of DDIM sampling are required, involving 25 noise predictions by the U-Net.
+The denoising process removes noise at timestamp $t$ using a sampling recurrence formula to obtain the vector $z_{t-1}$, gradually leading $z_{0}$, and finally uses an image decoder $\Psi$ to restore the image.
+During the gradient update process, each step of DDIM sampling can be expressed as:
 […]
 Each denoising step requires computing and storing the U-Net gradient, which can lead to memory overflow.
-However, since the predicted noise $\varepsilon_{\theta}(z_{t},t,\mathcal{C})$ follows a standard normal distribution $\mathcal{N}(0,I)$, we can use the reparameterization trick [10] to bypass the gradient computation of the U-Net and obtain an approximate gradient:
+To address this, we adopt the approach from previous work [11] by bypassing the gradient computation of the U-Net, thereby obtaining an approximate gradient.
+When optimizing the vector $z_{T/2}$ with the loss function $\mathcal{L}$, it can be expanded using the chain rule as follows:
 $\nabla_{z_{T/2}}\mathcal{L}\approx\frac{\partial\mathcal{L}}{\partial x}\cdot\frac{\partial x}{\partial z_{p}}\cdot\frac{\partial z_{p}}{\partial z_{0}}\cdot\sqrt{\frac{1}{\alpha_{1}}}\cdot\cdots\cdot\sqrt{\frac{\alpha_{T/2-1}}{\alpha_{T/2}}}.$ (9)
+where $z_{p}$ represents the adversarial patch obtained through the image encoder $\Psi$, i.e., $z_{p}=\Psi(z_{0})$ , and $x$ is the training data with $z_{p}$ attached.
 3.3 Target Mask Control
-Our method can generate adversarial patches of various shapes, with the process being controlled by the target mask $m$.
-Using target masks serves not only to extract patches of different shapes but also plays a crucial role in the natural appearance of the patch.
+Our method can generate adversarial patches of various shapes, controlled by the target mask $m$.
+Using target masks not only allows for the extraction of patches of different shapes but also plays a crucial role in maintaining the natural appearance of the patch.
 Since Stable Diffusion [45] generates square images, there is environmental information surrounding our target.
-After selecting the initial image $\mathcal{I}$, to prevent the background from interfering with the main subject during the optimization process, we need to replace the background with a solid-colored image $s$ before the Image-to-Latent Mapping stage:
-$\mathcal{I}^{\prime}=\mathcal{I}\odot m+s\odot(1-m),$ (11)
+After selecting the initial image $\mathcal{I}$, to prevent the background from interfering with the main subject during the optimization process, we replace the background with a solid-colored image $s$ before the Image-to-Latent Mapping stage:
+$\mathcal{I}^{\prime}=\mathcal{I}\odot m+s\odot(1-m).$ (10)
 Additionally, to ensure that the background remains unchanged throughout the update process, we downsample the mask $m$ by a factor of 8 to match the size of latent vectors and overlay gradients outside the target area.
-Overall, the DiffPatch generation process starts with Image-to-Latent Mapping.
-Incomplete Diffusion Optimization is then applied to reduce memory consumption and enhance performance.
-Finally, Target Mask Control is used to maintain a natural appearance.
-The generation process is illustrated in Algorithm 1.
-In addition, the patch can be iteratively optimized using DiffPatch to achieve an accumulative effect, as shown in Figure 3.
-Specifically, the generated patch can be used to generate a new patch, enabling stronger attack performance while maintaining natural appearance.
-Figure 3: The effect of iterative generation (accumulated style variation from left to right).
+Overall, the DiffPatch generation process begins with Image-to-Latent Mapping, followed by Incomplete Diffusion Optimization to reduce memory consumption and enhance performance.
+Finally, Target Mask Control is applied to preserve the natural appearance of the generated patches.
+The complete generation process is outlined in Algorithm 1.
+Furthermore, the patch can be iteratively optimized using DiffPatch to achieve an accumulative effect, as illustrated in Figure 3.
+Specifically, the generated patch can serve as the input for generating a new patch, enabling stronger attack performance while maintaining a natural appearance.
+This iterative approach ensures both effectiveness and stealthiness in adversarial patch generation.
+Figure 3: The effect of iterative generation (showing accumulated style variation from left to right).
 Algorithm 1 DiffPatch Generation Process
@@
 4 Experiments
-[…]
-Method YLv3 YLv3t YLv4 YLv4t YLv5s YLv7t FRCNN DETR
-AP |||| ASR AP |||| ASR AP |||| ASR AP |||| ASR AP |||| ASR AP |||| ASR AP |||| ASR AP |||| ASR
-${}^{(P_{A})}$dvYL 31.9 |||| 46.3 13.6 |||| 66.4 46.3 |||| 36.2 9.0 |||| 69.0 24.8 |||| 53.4 61.8 |||| 20.7 38.3 |||| 24.0 33.1 |||| 36.7
-${}^{(P_{B})}$-SEA 22.0 |||| 39.0 7.9 |||| 59.0 18.4 |||| 48.7 12.2 |||| 56.8 10.2 |||| 49.3 46.3 |||| 26.6 20.0 |||| 27.1 21.8 |||| 36.3
-${}^{(P_{C})}$ 50.4 |||| 33.7 15.9 |||| 74.2 61.3 |||| 26.2 13.7 |||| 78.9 41.0 |||| 41.4 69.0 |||| 17.5 55.5 |||| 25.1 43.3 |||| 41.5
-${}^{(P_{D})}$2D 50.3 |||| 33.9 18.8 |||| 69.1 58.4 |||| 30.9 15.6 |||| 75.0 41.0 |||| 40.7 68.7 |||| 18.4 52.9 |||| 25.0 44.8 |||| 34.2
-${}^{(P_{1})}$3(Ours) 12.9 |||| 70.4 17.0 |||| 64.1 58.2 |||| 28.1 21.3 |||| 60.4 36.0 |||| 52.2 60.2 |||| 24.6 38.6 |||| 29.3 38.6 |||| 38.6
-${}^{(P_{2})}$3t(Ours) 40.2 |||| 38.6 3.9 |||| 86.1 62.5 |||| 24.4 16.9 |||| 65.3 40.6 |||| 44.8 59.4 |||| 24.1 44.1 |||| 29.5 40.4 |||| 39.6
-${}^{(P_{3})}$4(Ours) 32.3 |||| 45.0 23.8 |||| 57.5 35.1 |||| 49.9 31.1 |||| 46.8 45.1 |||| 43.5 62.4 |||| 21.6 43.7 |||| 27.4 39.0 |||| 39.0
-${}^{(P_{4})}$4t(Ours) 40.8 |||| 41.0 11.8 |||| 73.6 61.6 |||| 25.5 9.2 |||| 79.1 38.3 |||| 50.3 57.4 |||| 25.7 45.3 |||| 28.4 39.5 |||| 41.4
-${}^{(P_{5})}$5s(Ours) 33.9 |||| 47.4 17.8 |||| 64.3 60.3 |||| 27.2 23.7 |||| 56.5 22.0 |||| 69.9 61.3 |||| 23.1 40.6 |||| 29.9 37.2 |||| 38.3
-${}^{(P_{6})}$7t(Ours) 49.6 |||| 35.0 19.9 |||| 64.3 64.2 |||| 23.0 26.0 |||| 56.7 39.6 |||| 44.5 51.7 |||| 33.5 50.2 |||| 26.1 42.2 |||| 38.9
-${}^{(P_{7})}$(Ours) 38.8 |||| 42.7 19.9 |||| 62.4 58.1 |||| 29.2 23.0 |||| 60.8 37.2 |||| 50.7 61.7 |||| 24.3 22.6 |||| 38.5 35.4 |||| 43.7
-${}^{(P_{8})}$(Ours) 46.6 |||| 35.2 35.2 |||| 49.0 67.1 |||| 22.1 34.3 |||| 46.7 44.6 |||| 43.1 66.8 |||| 20.5 46.2 |||| 26.1 37.1 |||| 42.0
-${}^{(P_{j})}$ource 53.5 |||| 29.7 45.9 |||| 38.1 70.0 |||| 18.5 47.8 |||| 36.9 55.8 |||| 35.0 69.4 |||| 18.4 53.5 |||| 23.3 47.3 |||| 33.6
-       Gray 72.3 |||| 16.5 61.8 |||| 25.2 76.8 |||| 12.8 56.4 |||| 30.0 74.1 |||| 15.5 77.1 |||| 13.5 64.4 |||| 17.3 50.2 |||| 34.4
-${}^{(P_{m})}$andom 70.2 |||| 16.9 64.3 |||| 22.2 78.5 |||| 12.2 60.4 |||| 26.9 75.1 |||| 15.5 79.8 |||| 12.3 67.3 |||| 17.6 51.1 |||| 31.7
-${}^{(P_{n})}$andom 72.8 |||| 17.0 68.0 |||| 19.0 82.4 |||| 10.4 69.8 |||| 20.7 83.2 |||| 11.7 79.3 |||| 12.4 72.0 |||| 16.4 57.0 |||| 25.9
-$P_{1}$ $P_{2}$ $P_{3}$ $P_{4}$ $P_{5}$ $P_{6}$ $P_{7}$ $P_{8}$ $P_{A}$ $P_{B}$ $P_{C}$ $P_{D}$ $P_{j}$ $P_{m}$ $P_{n}$
+Table 1: Comparing DiffPatch with other methods in ASR (%) (higher is better).
+The best results are underlined.
+denotes transfer attack methods, while the rest are white-box attacks.
+trained on YLv2, trained on YLv4t.
+The generated patches are shown at the bottom.
+Method ASR ($\uparrow$)
+YLv3 YLv3t YLv4 YLv4t YLv5s YLv7t YLv10s YLv12s FRCNN DETR RT-DETR
+${}^{(P_{A})}$dvYL 46.3 66.4 36.2 69.0 53.4 20.7 35.1 30.6 24.0 36.7 30.6
+${}^{(P_{B})}$-SEA 39.0 59.0 48.7 56.8 49.3 26.6 61.1 62.1 27.1 36.3 32.1
+${}^{(P_{C})}$ 33.7 74.2 26.2 78.9 41.4 17.5 33.4 22.0 25.1 41.5 48.3
+${}^{(P_{D})}$2D 33.9 69.1 30.9 75.0 40.7 18.4 23.1 26.0 25.0 34.2 39.7
+${}^{(P_{E})}$ 39.0 49.5 32.9 46.0 43 18.3 20.7 23.0 22.9 39.8 25.6
+${}^{(P_{F})}$dvART 22.9 36.4 15.4 33.9 19.3 17.3 12.1 11.4 20.4 31.0 19.0
+${}^{(P_{1})}$v3(Ours) 70.4 64.1 28.1 60.4 52.2 24.6 26.6 26.8 29.3 38.6 44.9
+${}^{(P_{2})}$v3t(Ours) 38.6 86.1 24.4 65.3 44.8 24.1 24.0 19.9 29.5 39.6 46.7
+${}^{(P_{3})}$v4(Ours) 45.0 57.5 49.9 46.8 43.5 21.6 27.5 24.8 27.4 39.0 45.2
+${}^{(P_{4})}$v4t(Ours) 41.0 73.6 25.5 79.1 50.3 25.7 27.8 21.0 28.4 41.4 46.5
+${}^{(P_{5})}$v5s(Ours) 47.4 64.3 27.2 56.5 69.9 23.1 31.6 25.7 29.9 38.3 37.7
+${}^{(P_{6})}$v7t(Ours) 35.0 64.3 23.0 56.7 44.5 33.5 31.6 25.3 26.1 38.9 47.7
+${}^{(P_{7})}$v10s(Ours) 35.3 62.0 23.4 59.9 45.7 24.2 66.1 27.2 28.9 40.3 45.0
+${}^{(P_{8})}$v12s(Ours) 44.0 57.3 37.9 55.1 45.1 22.2 47.7 63.7 32.4 27.4 42.7
+${}^{(P_{9})}$(Ours) 42.7 62.4 29.2 60.8 50.7 24.3 28.8 24.8 38.5 43.7 45.5
+${}^{(P_{10})}$(Ours) 64.2 79.5 43.1 75.2 62.7 27.1 41.2 31.8 37.0 46.5 52.0
+${}^{(P_{11})}$-DETR(Ours) 36.9 55.2 22.2 48.5 40.1 22.3 26.7 23.5 27.1 37.9 56.3
+${}^{(P_{j})}$ource 29.7 38.1 18.5 36.9 35.0 18.4 16.0 17.4 23.3 33.6 27.3
+       Gray 16.5 25.2 12.8 30.0 15.5 13.5 13.2 12.3 17.3 34.4 16.5
+${}^{(P_{m})}$andom 16.9 22.2 12.2 26.9 15.5 12.3 11.2 12.7 17.6 31.7 15.6
+${}^{(P_{n})}$andom 17.0 19.0 10.4 20.7 11.7 12.0 9.8 10.6 16.4 25.9 15.5
+$P_{1}$      $P_{2}$      $P_{3}$      $P_{4}$      $P_{5}$      $P_{6}$      $P_{A}$      $P_{B}$      $P_{C}$      $P_{m}$
+[Uncaptioned image]      [Uncaptioned image]      [Uncaptioned image]      [Uncaptioned image]      [Uncaptioned image]      [Uncaptioned image]      [Uncaptioned image]      [Uncaptioned image]      [Uncaptioned image]      [Uncaptioned image]
+$P_{7}$      $P_{8}$      $P_{9}$      $P_{10}$      $P_{11}$      $P_{j}$      $P_{D}$      $P_{E}$      $P_{F}$      $P_{n}$
 4.1 Experimental Setup
 Datasets
-In our experiments, we trained and evaluated the adversarial patches using the INRIA person dataset [11], which consists of 614 training images and 288 test images.
-To meet the input requirements of the detector, we padded all images with gray pixels to make them square and then resized them to 640 $\times$ 640.
-Direct resizing would distort the image proportions.
-Furthermore, we evaluated the transferability of different adversarial patch methods on the MPII Human Pose dataset [1], COCO dataset [31] and 1,000 images from diverse scenarios collected from the Internet.
-Attacked Models
-Our experiments attack a variety of detectors, including YOLOv3, YOLOv3-tiny [37], YOLOv4, YOLOv4-tiny [4], YOLOv5s [25], YOLOv7-tiny [44], FasterRCNN [38] and DETR [6].
+In our experiments, we train and evaluate the adversarial patches on the INRIA person dataset [12], which consists of 614 training images and 288 test images.
+To meet the input requirements of the detector, we pad all images with gray pixels to make them square and then resize them to 640 $\times$ 640.
+Furthermore, we evaluate the transferability of different adversarial patch methods on the MPII Human Pose dataset [2], COCO dataset [35] and 1,000 images from diverse scenarios collected from the Internet.
+Victim Models
+Our target a variety of detectors, including YOLOv3, YOLOv3-tiny [43], YOLOv4, YOLOv4-tiny [5], YOLOv5s [28], YOLOv7-tiny [53], YOLOv10s [52], YOLOv12s [51], FasterRCNN [44] and DETR [7] and RT-DETR [60].
 We will abbreviate YOLO as YL and FasterRCNN as FRCNN for ease of presentation, such as YOLOv4-tiny as YLv4t.
-All models are pre-trained on COCO dataset [31].
+All models were pre-trained on the COCO dataset [35].
 Implementation Details
-We treat person as the target attack class, set the batch size to 32, the maximum epoch $N_{iter}=200$, DDIM sampling steps $T=50$, unconditional embedding optimization iterations $N_{u}=10$, $\eta_{u}=0.01$, $\eta_{p}=0.003$, $\epsilon=0.5$, we use Adam optimizer [26] with the default parameters, the textual prompts $\mathcal{P}$ are automatically generated using BLIP [29], the version of Stable Diffusion [39] is v1.4.
-Since we are generating irregular patches, we set the patch size to match the pixel area of a square patch with a scale of $\tau=0.2$.
-To achieve stronger performance in patch attacks, we applied two iterations of optimization to the patch using DiffPatch.
-We regard the prediction of detector on clean data as the ground truth, and measure the capability of adversarial patch attack using the Average Precision (AP).
-A lower AP indicates that the accuracy of the detector in recognizing persons is reduced, indirectly indicating stronger patch attack performance.
-Additionally, since AP only measures the overall decline trend caused by the patches and does not directly reflect the attack performance, we also evaluated the Attack Success Rate (ASR) of various patches.
+We treat the person as the target object class, setting the batch size to 32, the maximum number of epochs $N_{iter}=200$, DDIM sampling steps $T=50$, unconditional embedding optimization iterations $N_{u}=10$, learning rates $\eta_{u}=0.01$ and $\eta_{p}=0.003$, and perturbation constraint $\epsilon=0.5$.
+We use the Adam optimizer [29] with its default parameters.
+Textual prompts $\mathcal{P}$ are automatically generated using BLIP [33].
+The adopted version of Stable Diffusion [45] is v1.4.
+Since we generate irregular patches, we set the patch size to match the pixel area of a square patch with a scale of $\tau=0.2$.
+To achieve stronger performance, we apply two iterations of optimization to the patch using DiffPatch.
+We evaluate the effectiveness of different patch generation methods using both the Attack Success Rate (ASR) and Average Precision (AP).
+The ASR measures the success of the patches in misleading the detector, while AP reflects the overall decline in detection accuracy.
 The confidence threshold and IoU threshold for the detector are both set to 0.5.
-All of our experiments were conducted on a single NVIDIA Tesla A100.
+All experiments were conducted on a single NVIDIA A100 GPU.
 4.2 Attack Performance Evaluation
-We generated a series of adversarial patches targeting different models using DiffPatch and compared them with current patch generation methods on INRIA dataset, including AdvYL [43], T-SEA [22], NPAP [21], and D2D [30].
-The detailed results can be seen in Table 1.
-DiffPatch ($P_{1}-P_{8}$) performs comparably to the best existing non-natural patch generation methods in terms of the AP metric.
-Notably, DiffPatch achieved the best Attack Success Rate (ASR), with an improvement of 10.1% over AdvYL on YLv4t.
-Compared to naturalistic adversarial patches ($P_{C}$ and $P_{D}$), our patch trained on YLv4t effectively reduced the AP by 4.5% and 6.4%, and increased the ASR by 0.2% and 4.1%.
+We generate a series of adversarial patches targeting different models using DiffPatch and compare them on the INRIA dataset with existing methods, including AdvYL [50], T-SEA [25], DAP [20], AdvART [19], NPAP [24], and D2D [34].
+The detailed results are reported in Table 1.
+As shown, DiffPatch ($P_{1}-P_{11}$) achieves the highest Attack Success Rate (ASR), improving by 10.1% over AdvYL on YOLOv4-tiny.
+When compared to naturalistic adversarial patches ($P_{C}$ and $P_{D}$), our patch trained on YOLOv4 results in a 0.2% and 4.1% ASR increase, respectively.
+Additionally, as demonstrated in the Appendix B, DiffPatch exhibits strong performance in the AP metric, performing comparably to unnatural patch generation methods.
 4.3 Cross-dataset Evaluation
-Considering the relatively simple scenes in the INRIA dataset, we conducted tests to evaluate the attack performance of the patches in more diverse environments.
-We trained patches using YLv4t with various methods and tested them on multiple datasets, including MPII and COCO.
-Due to the large size of these datasets, we randomly sampled a subset of images containing people for testing, selecting 729 images from COCO and 569 images from MPII.
-Additionally, since the images from these two datasets are relatively outdated with low-resolution images, we also collected 1,000 high-resolution images from the Internet, representing diverse scenes (denoted as “Collected”).
-The results, as shown in Table 2, demonstrate that our method significantly reduces the AP while substantially increasing the ASR.
+Given the relatively simple scenes in the INRIA dataset, we further evaluate the attack performance of the patches in more diverse environments.
+We train patches using YOLOv4 with various methods and test them across multiple datasets, including MPII and COCO.
+Due to the large size of these datasets, we randomly sample a subset of images containing people for testing, specifically 729 images from COCO and 569 images from MPII.
+Additionally, since the images from these two datasets are relatively outdated and of low resolution, we also collect 1,000 high-resolution images from the Internet, representing diverse scenes (denoted as ”Collected”).
+As shown in Table 2, our method significantly reduces the AP while substantially increasing the ASR.
 On average, the ASR shows a 17.5% improvement over the unnatural patch method T-SEA and a 2.6% improvement over the natural patch method NPAP.
+Several example images and their detection results with adversarial patches are provided in the Appendix A.
 Metric Dataset Attack
 T-SEA NPAP D2D Ours
+ASR%$\uparrow$ INRIA 56.8 78.9 75.0 79.1
+MPII 60.3 73.4 73.3 75.9
+COCO 54.5 63.9 60.3 67.8
+Collected 67.2 82.1 79.4 85.8
 AP%$\downarrow$ INRIA 12.2 13.7 15.6 9.2
@@
 Collected 8 6.7 7.2 3.9
-ASR%$\uparrow$ INRIA 56.8 78.9 75.0 79.1
-MPII 60.3 73.4 73.3 75.9
-COCO 54.5 63.9 60.3 67.8
-Collected 67.2 82.1 79.4 85.8
-Table 2: Attack performance of different attacks across various datasets.
-4.4 Ensemble Attack Evaluation
-We employed an ensemble approach, training our patch using the averaged gradients from multiple models.
-Compared to other natural patch generation methods, our single patch achieved strong attack performance across various object detection models.
-As shown in Table 3, NPAP was trained on YLv3 and YLv4t, while D2D and our DiffPatch were trained on all models.
-In white-box attacks, our method demonstrated a significant performance improvement over D2D.
+Table 2: Attack performance of different methods across datasets.
+Ensemble Attack
+We also evaluate an ensemble version of DiffPatch, which trains the patch using averaged gradients from multiple models.
+We compare this with two other methods, NPAP and D2D, where NPAP was trained on YOLOv3 and YOLOv4, and D2D was trained on all considered models.
+Our ensemble DiffPatch was similarly trained on all models.
+The results, presented in the Appendix C, demonstrate that our method outperforms NPAP and D2D by a significant margin.
 […]
-Method YLv3 YLv3t YLv4 YLv4t YLv5s YLv7t FRCNN Avg.
-AP |||| ASR AP |||| ASR AP |||| ASR AP |||| ASR AP |||| ASR AP |||| ASR AP |||| ASR AP |||| ASR
-NPAP 45.2 |||| 34.2 16.2 |||| 74.8 65.8 |||| 22.6 19.3 |||| 71.7 57.4 |||| 28.5 70.8 |||| 17.3 57.8 |||| 24.0 32.3 |||| 53.0
-D2D 48.9 |||| 35.7 17.5 |||| 66.8 58.4 |||| 28.9 24.6 |||| 59.3 35.6 |||| 43.2 67.0 |||| 19.5 52.7 |||| 24.3 43.5 |||| 39.7
-DiffPatch 9.1 |||| 65.9 6.1 |||| 79.5 28.9 |||| 47.4 12.0 |||| 75.5 23.2 |||| 62.4 55.4 |||| 28.5 19.9 |||| 38.9 22.1 |||| 56.9
+$Patch$
+20406080Average ASR(%)406080Average CLIP Sim.(%)ASRCLIP Sim.
+Figure 4: The effect of optimization constraint $\epsilon$ on latent vectors.
 4.4 Ablation Studies
+Optimization Constraint
+To ensure the natural appearance of the generated patches, we constrain each optimization step using the $L_{\infty}$ norm.
+The choice of $\epsilon$ will influence visual naturalness of the generated patch and its attack performance.
+To quantify the natural appearance of the patches, we use GPT-4o [1] to generate descriptive text for reference images (denoted by ’-’) and compute the CLIP similarity [41] between these descriptions and the generated patches.
+The corresponding descriptions can be found in the Appendix D.
+As illustrated in Figure 4, we obtain patches with varying styles under different constraints using DiffPatch and evaluate both the ASR and CLIP similarity for each.
+Based on the results, we select $\epsilon=0.5$ as the optimal value, as it achieves a favorable balance between attack performance and natural appearance.
 Loss Function
-We conducted experiments using both the Common Detection Loss and IoU-Detection Loss under the same patch training setting.
-As shown in Figure 4, compared to the Common Detection Loss, our IoU-Detection Loss demonstrates greater stability during training process and achieves higher attack performance in a earlier time.
+Here, we compare our IoU-Detection Loss with the Common Detection Loss under the same patch training setting.
+As shown in Figure 5, compared to the Common Detection Loss, our IoU-Detection Loss demonstrates greater stability during training process and achieves higher attack performance in an earlier time.
 Figure 5: Comparison of two loss functions. Left: training loss; Right: attack performance.
 Shape Impact
-The impact of patch shapes was investigated, as indicated by the data for $P_{m}$ and $P_{n}$ in Table 1.
-We compared the performance of random noise images with square and irregular shapes (with the same size as DiffPatch).
-From the AP and ASR metrics, it is evident that square patches yield higher attack benefits compared to irregular patches.
+Here, we investigate the impact of patch shapes, as used for $P_{m}$ and $P_{n}$ in Table 1.
+We compare the performance of random noise images with square and irregular shapes (of the same size as DiffPatch).
+From the ASR and AP results in Table 1, it is evident that square patches yield higher attack effectiveness compared to irregular patches.
 In other words, designing irregular patches with high attack success rates is more challenging.
 Diffusion Trajectory Length
-Using latent vectors from intermediate timestamps to generate patches effectively mitigates semantic loss.
-We tested the results of latent vectors at different timestamps, evaluating the attack performance and similarity to textual descriptions for patches generated from diffusion trajectories of varying lengths.
-See Figure 5 for details.
-We use the cosine similarity of CLIP [35] embeddings to measure the alignment between the patch and the prompt, which indirectly reflects the visual quality of the patch.
-It is evident that starting diffusion optimization from $\frac{T}{2}$ (25 Steps) not only achieves a high success rate but also produces images that better align with textual descriptions, thereby maintaining greater semantic consistency with the original images.
+Here, we evaluate the results of latent vectors at different timestamps, assessing both the attack performance and the similarity to textual descriptions for patches generated from diffusion trajectories of varying lengths.
+As shown in Figure 6, starting diffusion optimization from $\frac{T}{2}$ (25 steps) not only achieves a high success rate but also produces images that align more closely with textual descriptions, thereby maintaining greater semantic consistency with the original images.
 Figure 6: The effects of diffusion trajectories with different lengths on the performance of patch attacks and natural appearance.
@@
 The target mask mechanism helps mitigate semantic loss in images to some extent.
-We compared the training results with and without the use of this mechanism.
-Without using target mask leads to quality issues in the patches during optimization, as shown in Figure 6.
-Figure 6: The impact of target mask on adversarial patch optimization. Without target mask control, adversarial patches exhibit noticeable detail distortions (indicated by red boxes).
-Optimization Constraint
-We use the infinity norm to constrain each optimization of the patch to ensure its natural appearance.
-The size of $\epsilon$ significantly impacts patch generation.
-An appropriate $\epsilon$ can effectively balance the trade-off between attack performance and appearance.
-As shown in Table 4, we ultimately chose $\epsilon=0.5$ as the constraint parameter for patch generation, as it maintains both appearance and a high attack success rate.
-Constraint $\epsilon$ 0.3 0.5 0.7
-[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-CLIP Sim. 75.6% 73.2% 63.9%
-ASR 57.0% 63.0% 67.4%
-Table 4: The effect of optimization constraint $\epsilon$ on latent vectors.
+Here, we compare the training results with and without the use of this mechanism.
+Without the target mask, the patches exhibit quality issues during optimization, as illustrated in Figure 7.
 Iterative Optimization
-As the results in Table 5, iterative optimization can further enhance the adversarial strength of the patch. However, it also makes the patch unnatural.
+As shown in Table 3, iterative optimization can further enhance the adversarial strength of the patch. However, it also tends to reduce the natural appearance of the patch.
 Metric Iterative Model
 YLv3t YLv4t YLv5s FRCNN
+ASR%$\uparrow$ w/o 82.9 75.5 65.6 37.3
+w/ 86.1 79.1 69.9 38.5
 AP%$\downarrow$ w/o 7.1 12.3 24.6 26.0
 w/ 3.9 9.2 22.0 22.6
-ASR%$\uparrow$ w/o 82.9 75.5 65.6 37.3
-w/ 86.1 79.1 69.9 38.5
 Sim.$\uparrow$ w/o 77.1 77.6 73.2 72.9
@@
 Table 3: The impact of using iterative optimization.
-4.6 Physical-world Adversarial Patch Dataset
-Our DiffPatch achieves promising attack results even in physical-world environments.
-Using DiffPatch, we designed 9 adversarial patches and printed them on T-shirts, as shown in Figure 7.
-We recruited 20 participants (with ethics approval) to capture images in diverse indoor and outdoor, including laboratories, campus, cafeteria, subway station, and shopping mall.
-In total, we collected 1,131 images into our AdvPatch-1K dataset, including both individual and group (2 to 10+ persons) photos.
-The dataset is accompanied with detailed annotations of the person and patch locations (bounding boxes).
-Several example images are shown in Figure 1, with detailed statistics in Figure 8.
-The images in AdvPatch-1K can effectively evade the YOLOv5s detector, as shown in the right subfigure of Figure 8.
-Our dataset provides a valuable resource for advancing defense research against adversarial patch attacks.
-Figure 7: Adversarial T-shirts created using DiffPatch.
-Figure 8: Left: The proportion of images with different numbers of persons in AdvPatch-1K. Right: The detection performance of YOLOv5s on AdvPatch-1K under different confidence thresholds.
+4.5 Adversarial T-shirt Dataset
+To better evaluate the effectiveness of our attack and create a valuable physical-world adversarial patch dataset for future research, we designed 9 physical-world adversarial T-shirts using our generated patches.
+We printed these patches on T-shirts and captured photos in a variety of indoor and outdoor scenarios, such as laboratory, campus, cafeteria, subway station, and shopping mall, with the assistance of 20 participants (with ethics approval).
+In total, we collected 1,131 images, including both individual and group photos (ranging from 2 to 10+ persons).
+We name the dataset AdvT-shirt-1K, which includes detailed annotations for person and patch locations (bounding boxes).
+The collection and annotation process of the dataset spanned six months, which involves the refinement of our DiffPatch method for multiple times.
+The statistics of the dataset is shown in Figure 8 with several example images are provided in Figure 1.
+The results shown in the right subfigure of Figure 8 demonstrate the effectiveness of the adversarial T-shirts in evading the YOLOv5s detector.
+Although small in scale, our dataset provides a practical resource for developing and testing the real-world effectiveness of adversarial patch detectors.
+Figure 7: The impact of target mask on adversarial patch optimization. Without target mask control, adversarial patches exhibit noticeable detail distortions (indicated by the red boxes).
+Figure 8: Left: The distribution of the images in AdvT-shirt-1K. Right: The detection performance of YOLOv5s on AdvT-shirt-1K under different confidence thresholds.
 5 Conclusion
@@
 Extensive experiments across multiple object detection models validate the effectiveness of DiffPatch.
-With DiffPatch, we also created a physical-world adversarial dataset, AdvPatch-1K, comprising 1,131 images captured in diverse scenes and conditions.
-AdvPatch-1K can help develop robust defense methods against adversarial patch attacks.
+With DiffPatch, we also created a physical-world adversarial dataset, AdvT-shirt-1K, comprising 1,131 images captured in diverse scenes and conditions, which validates the feasibility of DiffPatch in physical-world environments.

### v5

@@
-DiffPatch: Generating Customizable Adversarial Patches using Diffusion Models
+BadPatch: Diffusion-Based Generation of Physical Adversarial Patches
 Abstract
@@
 While generative adversarial networks and diffusion models can produce more natural-looking patches, they often fail to balance stealthiness with attack effectiveness and lack flexibility for user customization.
-To address these limitations, we propose DiffPatch, a novel diffusion-based framework for generating customizable and naturalistic adversarial patches.
+To address these limitations, we propose BadPatch, a novel diffusion-based framework for generating customizable and naturalistic adversarial patches.
 Our approach allows users to start from a reference image (rather than random noise) and incorporates masks to create patches of various shapes, not limited to squares.
@@
 Our method achieves attack performance comparable to state-of-the-art non-naturalistic patches while maintaining a natural appearance.
-Using DiffPatch, we construct AdvT-shirt-1K, the first physical adversarial T-shirt dataset comprising over a thousand images captured in diverse scenarios.
+Using BadPatch, we construct AdvT-shirt-1K, the first physical adversarial T-shirt dataset comprising over a thousand images captured in diverse scenarios.
 AdvT-shirt-1K can serve as a useful dataset for training or testing future defense methods.
+The code and datasets are available at https://github.com/Wwangb/BadPatch.
 1 Introduction
-Figure 1: Example images from our AdvT-shirt dataset, showcasing the performance of DiffPatch in various scenarios.
+Figure 1: Example images from our AdvT-shirt dataset, showcasing the performance of BadPatch in various scenarios.
 Object detectors powered by deep neural networks (DNNs) are integral to modern computer vision systems, enabling critical applications such as autonomous driving [9] and medical diagnostics [40].
-However, DNNs are highly susceptible to adversarial attacks [8, 57, 16, 55], which manipulate their outputs in ways that defy logic or compromise safety.
+However, DNNs are highly susceptible to adversarial attacks [8, 59, 16, 57, 38, 55], which manipulate their outputs in ways that defy logic or compromise safety.
 These attacks can cause object detectors to misclassify or fail to detect critical objects.
@@
 For instance, recent work [11] leverages image-latent mapping for adversarial sample generation, but its large perturbations in physical settings lead to poor naturalness or effectiveness.
-To address these limitations, we propose DiffPatch, a novel adversarial patch generation framework based on Stable Diffusion [45].
-DiffPatch generates naturalistic and customizable adversarial patches from a reference image specified by the adversary.
+To address these limitations, we propose BadPatch, a novel adversarial patch generation framework based on Stable Diffusion [46].
+BadPatch generates naturalistic and customizable adversarial patches from a reference image specified by the adversary.
 Our approach first maps the reference image into the latent space and optimizes noise latent vectors to create adversarial patches.
@@
 To generate irregular-shaped patches, we use masks to replace the background and suppress gradients in non-essential areas, reducing interference during optimization.
-Using DiffPatch, we construct AdvT-shirt-1K, a dataset of over a thousand images featuring customizable adversarial patches in diverse physical-world scenarios.
+Using BadPatch, we construct AdvT-shirt-1K, a dataset of over a thousand images featuring customizable adversarial patches in diverse physical-world scenarios.
 This dataset demonstrates the practicality of our approach and provides a valuable resource for future research.
 In summary, our key contributions are as follows:
-  • We propose a novel diffusion-based method called DiffPatch for generating naturalistic adversarial patches based on a reference image.
-DiffPatch integrates Null-text inversion, Incomplete Diffusion Optimization (IDO) , masked customization, and an IoU-Detection loss to generate stylized and customizable adversarial patches.
-  • Experimental results demonstrate that our DiffPatch achieves the highest attack success rate (ASR), significantly reducing the mean average precision (mAP) of various object detectors.
+  • We propose a novel diffusion-based method called BadPatch for generating naturalistic adversarial patches based on a reference image.
+BadPatch integrates Null-text inversion, Incomplete Diffusion Optimization (IDO) , masked customization, and an IoU-Detection loss to generate stylized and customizable adversarial patches.
+  • Experimental results demonstrate that our BadPatch achieves the highest attack success rate (ASR), significantly reducing the mean average precision (mAP) of various object detectors.
 Its performance rivals or even surpasses state-of-the-art methods for generating unnatural adversarial patches.
@@
 These advancements highlight the growing potential of generative models in crafting stealthy and robust adversarial attacks.
-Figure 2: An overview of the proposed DiffPatch generation framework.
+Figure 2: An overview of the proposed BadPatch generation framework.
 2.3 Diffusion-based Image Editing
@@
 Overview
-Figure 2 illustrates the overall framework of DiffPatch.
+Figure 2 illustrates the overall framework of BadPatch.
 It optimizes Null-text embeddings along incomplete diffusion trajectories and maps the image to latent vectors at intermediate timestamps, ensuring that the patch retains a natural appearance during updates.
@@
 To address memory constraints, we skip the U-Net gradients during optimization and use target masks to eliminate semantic interference from irrelevant regions, enabling the generation of more naturalistic adversarial patches.
-Next, we will introduce the key techniques of DiffPatch.
+Next, we will introduce the key techniques of BadPatch.
 3.1 Image-to-Latent Mapping
@@
 $\tilde{\varepsilon}_{\theta}(z_{t},t,\mathcal{C},\phi)=w\cdot\varepsilon_{\theta}(z_{t},t,\mathcal{C})+(1-w)\ \cdot\varepsilon_{\theta}(z_{t},t,\phi),$ (2)
-where $w$ is the guidance scale parameter, and $\phi=\mathcal{E}("")$ is the embedding of a null text.
-The generated content is highly sensitive to textual prompts, and even minor variations can cause trajectory deviations during guided generation.
-These deviations accumulate, making reconstruction challenging, especially during image editing.
+[…]
 To address this, we separately optimize the null text embedding $\phi_{t}$ for each timestamp $t$ in the diffusion process $t=T\rightarrow t=1$, guiding the process back to the correct trajectory.
@@
 Additionally, to ensure that the background remains unchanged throughout the update process, we downsample the mask $m$ by a factor of 8 to match the size of latent vectors and overlay gradients outside the target area.
-Overall, the DiffPatch generation process begins with Image-to-Latent Mapping, followed by Incomplete Diffusion Optimization to reduce memory consumption and enhance performance.
+Overall, the BadPatch generation process begins with Image-to-Latent Mapping, followed by Incomplete Diffusion Optimization to reduce memory consumption and enhance performance.
 Finally, Target Mask Control is applied to preserve the natural appearance of the generated patches.
 The complete generation process is outlined in Algorithm 1.
-Furthermore, the patch can be iteratively optimized using DiffPatch to achieve an accumulative effect, as illustrated in Figure 3.
+Furthermore, the patch can be iteratively optimized using BadPatch to achieve an accumulative effect, as illustrated in Figure 3.
 Specifically, the generated patch can serve as the input for generating a new patch, enabling stronger attack performance while maintaining a natural appearance.
@@
 Figure 3: The effect of iterative generation (showing accumulated style variation from left to right).
-Algorithm 1 DiffPatch Generation Process
+Algorithm 1 BadPatch Generation Process
 Input: A patch image $\mathcal{I}^{\prime}$ with target mask $m$, prompt embedding $\mathcal{C}=\mathcal{E}(\mathcal{P})$, an object detector $\mathcal{D}$, training dataset $S$ and ground truth box $y$.
@@
 8:end for
-9:  $m^{*}\leftarrow Downsample(m)$
+[…]
 10:for $k=1,\ldots,N_{iter}$ do
@@
 19:end for
-20:  $z_{p}\leftarrow G(\bar{z}_{\frac{T}{2}}+\delta_{B},\frac{T}{2},\mathcal{C},\{\phi_{t}\}^{\frac{T}{2}}_{t=1})\odot m$
+[…]
 21:return $z_{p}$
 4 Experiments
-Table 1: Comparing DiffPatch with other methods in ASR (%) (higher is better).
+Table 1: Comparing BadPatch with other methods in ASR (%) (higher is better).
 The best results are underlined.
@@
 The generated patches are shown at the bottom.
-Method ASR ($\uparrow$)
+Method ASR (% $\uparrow$)
 YLv3 YLv3t YLv4 YLv4t YLv5s YLv7t YLv10s YLv12s FRCNN DETR RT-DETR
-${}^{(P_{A})}$dvYL 46.3 66.4 36.2 69.0 53.4 20.7 35.1 30.6 24.0 36.7 30.6
-${}^{(P_{B})}$-SEA 39.0 59.0 48.7 56.8 49.3 26.6 61.1 62.1 27.1 36.3 32.1
-${}^{(P_{C})}$ 33.7 74.2 26.2 78.9 41.4 17.5 33.4 22.0 25.1 41.5 48.3
-${}^{(P_{D})}$2D 33.9 69.1 30.9 75.0 40.7 18.4 23.1 26.0 25.0 34.2 39.7
-${}^{(P_{E})}$ 39.0 49.5 32.9 46.0 43 18.3 20.7 23.0 22.9 39.8 25.6
-${}^{(P_{F})}$dvART 22.9 36.4 15.4 33.9 19.3 17.3 12.1 11.4 20.4 31.0 19.0
-${}^{(P_{1})}$v3(Ours) 70.4 64.1 28.1 60.4 52.2 24.6 26.6 26.8 29.3 38.6 44.9
-${}^{(P_{2})}$v3t(Ours) 38.6 86.1 24.4 65.3 44.8 24.1 24.0 19.9 29.5 39.6 46.7
-${}^{(P_{3})}$v4(Ours) 45.0 57.5 49.9 46.8 43.5 21.6 27.5 24.8 27.4 39.0 45.2
-${}^{(P_{4})}$v4t(Ours) 41.0 73.6 25.5 79.1 50.3 25.7 27.8 21.0 28.4 41.4 46.5
-${}^{(P_{5})}$v5s(Ours) 47.4 64.3 27.2 56.5 69.9 23.1 31.6 25.7 29.9 38.3 37.7
-${}^{(P_{6})}$v7t(Ours) 35.0 64.3 23.0 56.7 44.5 33.5 31.6 25.3 26.1 38.9 47.7
-${}^{(P_{7})}$v10s(Ours) 35.3 62.0 23.4 59.9 45.7 24.2 66.1 27.2 28.9 40.3 45.0
-${}^{(P_{8})}$v12s(Ours) 44.0 57.3 37.9 55.1 45.1 22.2 47.7 63.7 32.4 27.4 42.7
-${}^{(P_{9})}$(Ours) 42.7 62.4 29.2 60.8 50.7 24.3 28.8 24.8 38.5 43.7 45.5
-${}^{(P_{10})}$(Ours) 64.2 79.5 43.1 75.2 62.7 27.1 41.2 31.8 37.0 46.5 52.0
-${}^{(P_{11})}$-DETR(Ours) 36.9 55.2 22.2 48.5 40.1 22.3 26.7 23.5 27.1 37.9 56.3
-${}^{(P_{j})}$ource 29.7 38.1 18.5 36.9 35.0 18.4 16.0 17.4 23.3 33.6 27.3
+$^{(P_{A})}$AdvYL 46.3 66.4 36.2 69.0 53.4 20.7 35.1 30.6 24.0 36.7 30.6
+$^{(P_{B})}$T-SEA 39.0 59.0 48.7 56.8 49.3 26.6 61.1 62.1 27.1 36.3 32.1
+$^{(P_{C})}$NPAP 33.7 74.2 26.2 78.9 41.4 17.5 33.4 22.0 25.1 41.5 48.3
+$^{(P_{D})}$D2D 33.9 69.1 30.9 75.0 40.7 18.4 23.1 26.0 25.0 34.2 39.7
+$^{(P_{E})}$DAP 39.0 49.5 32.9 46.0 43 18.3 20.7 23.0 22.9 39.8 25.6
+$^{(P_{F})}$AdvART 22.9 36.4 15.4 33.9 19.3 17.3 12.1 11.4 20.4 31.0 19.0
+$^{(P_{1})}$YLv3(Ours) 70.4 64.1 28.1 60.4 52.2 24.6 26.6 26.8 29.3 38.6 44.9
+$^{(P_{2})}$YLv3t(Ours) 38.6 86.1 24.4 65.3 44.8 24.1 24.0 19.9 29.5 39.6 46.7
+$^{(P_{3})}$YLv4(Ours) 45.0 57.5 49.9 46.8 43.5 21.6 27.5 24.8 27.4 39.0 45.2
+$^{(P_{4})}$YLv4t(Ours) 41.0 73.6 25.5 79.1 50.3 25.7 27.8 21.0 28.4 41.4 46.5
+$^{(P_{5})}$YLv5s(Ours) 47.4 64.3 27.2 56.5 69.9 23.1 31.6 25.7 29.9 38.3 37.7
+$^{(P_{6})}$YLv7t(Ours) 35.0 64.3 23.0 56.7 44.5 33.5 31.6 25.3 26.1 38.9 47.7
+$^{(P_{7})}$YLv10s(Ours) 35.3 62.0 23.4 59.9 45.7 24.2 66.1 27.2 28.9 40.3 45.0
+$^{(P_{8})}$YLv12s(Ours) 44.0 57.3 37.9 55.1 45.1 22.2 47.7 63.7 32.4 27.4 42.7
+$^{(P_{9})}$FRCNN(Ours) 42.7 62.4 29.2 60.8 50.7 24.3 28.8 24.8 38.5 43.7 45.5
+$^{(P_{10})}$DETR(Ours) 64.2 79.5 43.1 75.2 62.7 27.1 41.2 31.8 37.0 46.5 52.0
+$^{(P_{11})}$RT-DETR(Ours) 36.9 55.2 22.2 48.5 40.1 22.3 26.7 23.5 27.1 37.9 56.3
+$^{(P_{j})}$Source 29.7 38.1 18.5 36.9 35.0 18.4 16.0 17.4 23.3 33.6 27.3
     Gray 16.5 25.2 12.8 30.0 15.5 13.5 13.2 12.3 17.3 34.4 16.5
-${}^{(P_{m})}$andom 16.9 22.2 12.2 26.9 15.5 12.3 11.2 12.7 17.6 31.7 15.6
-${}^{(P_{n})}$andom 17.0 19.0 10.4 20.7 11.7 12.0 9.8 10.6 16.4 25.9 15.5
+$^{(P_{m})}$Random 16.9 22.2 12.2 26.9 15.5 12.3 11.2 12.7 17.6 31.7 15.6
+$^{(P_{n})}$Random 17.0 19.0 10.4 20.7 11.7 12.0 9.8 10.6 16.4 25.9 15.5
 $P_{1}$    $P_{2}$    $P_{3}$    $P_{4}$    $P_{5}$    $P_{6}$    $P_{A}$    $P_{B}$    $P_{C}$    $P_{m}$
@@
 Since we generate irregular patches, we set the patch size to match the pixel area of a square patch with a scale of $\tau=0.2$.
-To achieve stronger performance, we apply two iterations of optimization to the patch using DiffPatch.
+To achieve stronger performance, we apply two iterations of optimization to the patch using BadPatch.
 We evaluate the effectiveness of different patch generation methods using both the Attack Success Rate (ASR) and Average Precision (AP).
@@
 4.2 Attack Performance Evaluation
-We generate a series of adversarial patches targeting different models using DiffPatch and compare them on the INRIA dataset with existing methods, including AdvYL [50], T-SEA [25], DAP [20], AdvART [19], NPAP [24], and D2D [34].
+We generate a series of adversarial patches targeting different models using BadPatch and compare them on the INRIA dataset with existing methods, including AdvYL [51], T-SEA [25], DAP [20], AdvART [19], NPAP [24], and D2D [34].
 The detailed results are reported in Table 1.
-As shown, DiffPatch ($P_{1}-P_{11}$) achieves the highest Attack Success Rate (ASR), improving by 10.1% over AdvYL on YOLOv4-tiny.
+As shown, BadPatch ($P_{1}-P_{11}$) achieves the highest Attack Success Rate (ASR), improving by 10.1% over AdvYL on YOLOv4-tiny.
 When compared to naturalistic adversarial patches ($P_{C}$ and $P_{D}$), our patch trained on YOLOv4 results in a 0.2% and 4.1% ASR increase, respectively.
-Additionally, as demonstrated in the Appendix B, DiffPatch exhibits strong performance in the AP metric, performing comparably to unnatural patch generation methods.
+Additionally, as demonstrated in the Appendix B, BadPatch exhibits strong performance in the AP metric, performing comparably to unnatural patch generation methods.
 4.3 Cross-dataset Evaluation
@@
 T-SEA NPAP D2D Ours
-ASR%$\uparrow$ INRIA 56.8 78.9 75.0 79.1
+[…]
 MPII 60.3 73.4 73.3 75.9
@@
 Collected 67.2 82.1 79.4 85.8
-AP%$\downarrow$ INRIA 12.2 13.7 15.6 9.2
+[…]
 MPII 17.0 17.3 17.9 12.5
@@
 Ensemble Attack
-We also evaluate an ensemble version of DiffPatch, which trains the patch using averaged gradients from multiple models.
+We also evaluate an ensemble version of BadPatch, which trains the patch using averaged gradients from multiple models.
 We compare this with two other methods, NPAP and D2D, where NPAP was trained on YOLOv3 and YOLOv4, and D2D was trained on all considered models.
-Our ensemble DiffPatch was similarly trained on all models.
+Our ensemble BadPatch was similarly trained on all models.
 The results, presented in the Appendix C, demonstrate that our method outperforms NPAP and D2D by a significant margin.
 […]
-$Patch$
+PatchPatch
 20406080Average ASR(%)406080Average CLIP Sim.(%)ASRCLIP Sim.
@@
 The corresponding descriptions can be found in the Appendix D.
-As illustrated in Figure 4, we obtain patches with varying styles under different constraints using DiffPatch and evaluate both the ASR and CLIP similarity for each.
+As illustrated in Figure 4, we obtain patches with varying styles under different constraints using BadPatch and evaluate both the ASR and CLIP similarity for each.
 Based on the results, we select $\epsilon=0.5$ as the optimal value, as it achieves a favorable balance between attack performance and natural appearance.
@@
 Here, we investigate the impact of patch shapes, as used for $P_{m}$ and $P_{n}$ in Table 1.
-We compare the performance of random noise images with square and irregular shapes (of the same size as DiffPatch).
+We compare the performance of random noise images with square and irregular shapes (of the same size as BadPatch).
 From the ASR and AP results in Table 1, it is evident that square patches yield higher attack effectiveness compared to irregular patches.
@@
 YLv3t YLv4t YLv5s FRCNN
-ASR%$\uparrow$ w/o 82.9 75.5 65.6 37.3
+[…]
 w/ 86.1 79.1 69.9 38.5
-AP%$\downarrow$ w/o 7.1 12.3 24.6 26.0
+[…]
 w/ 3.9 9.2 22.0 22.6
-Sim.$\uparrow$ w/o 77.1 77.6 73.2 72.9
+[…]
 w/ 72.1 69.4 60.0 60.8
@@
 We name the dataset AdvT-shirt-1K, which includes detailed annotations for person and patch locations (bounding boxes).
-The collection and annotation process of the dataset spanned six months, which involves the refinement of our DiffPatch method for multiple times.
+The collection and annotation process of the dataset spanned six months, which involves the refinement of our BadPatch method for multiple times.
 The statistics of the dataset is shown in Figure 8 with several example images are provided in Figure 1.
@@
 5 Conclusion
-In this work, we proposed a novel diffusion-based adversarial patch generation method DiffPatch to generate naturalistic and customized patches based on a reference image.
-DiffPatch utilizes image-to-latent mapping, incomplete diffusion optimization, and target mask control to create stylized adversarial patches.
-Extensive experiments across multiple object detection models validate the effectiveness of DiffPatch.
-With DiffPatch, we also created a physical-world adversarial dataset, AdvT-shirt-1K, comprising 1,131 images captured in diverse scenes and conditions, which validates the feasibility of DiffPatch in physical-world environments.
+In this work, we proposed a novel diffusion-based adversarial patch generation method BadPatch to generate naturalistic and customized patches based on a reference image.
+BadPatch utilizes image-to-latent mapping, incomplete diffusion optimization, and target mask control to create stylized adversarial patches.
+Extensive experiments across multiple object detection models validate the effectiveness of BadPatch.
+With BadPatch, we also created a physical-world adversarial dataset, AdvT-shirt-1K, comprising 1,131 images captured in diverse scenes and conditions, which validates the feasibility of BadPatch in physical-world environments.